A Bibliography of Publications in the Journal of Discrete Mathematical Sciences and Cryptography

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

31 October 2012 Version 0.02

Title word cross-reference

# [438, 446, 452, 460, 467, 477].

(0, n) [120]. (1, m, n)_{2} [74, 121]. (1 − u_{2}) [339]. (2q + 1) [380]. (n_{3}) [16]. (q + 1)^{2} [143]. (S, T) [359]. (W_{n}: 2) [374]. 0 [130]. 2 [19, 144, 455]. 2^{31} − 69 [411]. 3 [160, 173, 427]. 3 − (2m, 2m|1, 2m|2|1) [299]. 4 [86]. 64 [454]. [. .] [464]. [0, 1, m, n]_{1} [143]. [0, 2, 3, 4, 5] [13]. [m, n]1k [335]. A [360]. AG(2, 7) [13]. α [64]. b [132, 156]. C_{5} × C_{5} [40]. C_{λ} [14]. C_{pk} [122]. d [462]. δ [150]. E2(v − w, 4, 1) [331]. ϵ [237]. F [64]. F_{λ} [14]. F_{n} [273]. G(xyz) [429]. G − +− [451]. γ [109]. γ_{bi} [109]. GF(N_{F} (√)) [117]. G × C_{2n} [86]. h/v [99]. H_{v} [95].

IF 2[u]/u_{2} − 1 [333]. IFp + uIFp + u_{2}IFp [339]. K [87, 132, 163, 173, 177, 186, 200, 201, 223, 230, 231, 261, 282, 419, 462]. K1, 3 [427]. K1, 7 [427]. K2n, 2n|nC4 [403]. l [60]. m [154, 155, 172]. m_{2}(h) [12]. mΘ [473]. n [168, 170, 205, 223, 273]. ω [100]. p [177, 395]. P(v, 4, e) [331]. P_{3} [113]. PG(2, 2) [257]. PG(2, 3) [142, 257, 304, 335]. PG(2s + 1, q) [67]. PG(3, q) [9, 74, 294]. PG(4, q) [143]. PG(r, q) [380, 419]. P Q [95]. q [217, 218, 383]. R(1, m) [299]. PG(2, 2) [116]. PG(2, 3) [179]. PG(3, q) [121]. s [60]. SFLASH^{v3} [133]. σ_{F} (O_{F}/√) [117]. t [168, 175, 205, 328, 388]. t^{n} [161]. v [100].

xyz = − + + [429]. yL/t [344]. Z(n) [432]. Z_{2} [414].

-algebras [173]. -algorithm [237]. -arcs [419]. -ary [132, 155, 172, 175, 177, 454]. -binomial [217, 218]. -caps [380, 419]. -center [186]. -chromatic [132, 156]. -codes [150]. -cosets [64]. -critical [462]. -cyclic [339]. -designs [113]. -distributive [130]. -edge [462]. -Enumeration [383]. -error [177, 282]. -excellent [109]. -flexible [109]. -fold [60]. -free [427]. -Fuzzy [395]. -generated [455]. -hypergroups [100]. -hyperideals [200, 201]. -hyperoperations [95]. -ideals [87]. -Implication-based [388]. -involution [230, 231]. -languages [150]. -minimally [19]. -norms [359]. -ordered [261]. [?]out-of-Chang:2006:AFE,Chang:2007:RNI. -pebbling [328]. -polygroups [64]. -power [170]. -regular [86, 163]. -secure [144]. -sets [143, 335]. -SFP [144]. -structures [95, 99]. -th [154]. -valent [464]. -vertex [344, 427].

03 [404]. 04 [413]. 05 [422]. 06 [431].

1 [198].

2 [81]. 20-sets [13]. 2010 [404, 413, 422, 431]. 2011 [438, 446, 452, 460, 467, 477].

3-coloring [57]. 32-bit [188, 420]. 320 [297]. 320-bit [297].

4217th [411]. 4217th-order [411].

5 [121].

A. [260]. absolutely [179]. abstract [316]. accelerate [237]. access [22, 111, 158, 196, 286, 447]. accurate [385]. actions [292]. acyclic [115]. Ad [289]. adaptive [313]. Addendum [121]. addressing [128]. adjacency [274]. administration [423]. admission [341, 371]. AEA [63]. Affine [120, 263]. agent [367]. agents [367]. agglomerative [474]. agreement [239, 364, 469]. AHP [268]. Algebraic [20, 195, 234, 352, 473, 479]. algebras [20, 173]. Algorithm [15, 47, 72, 107, 164, 167, 186, 222, 223, 233, 237, 241, 243, 279, 282, 302, 306, 313, 360, 369, 393, 428, 433, 448, 466, 474]. algorithms [70, 71, 77, 103, 112, 127, 140, 154, 267, 358, 409]. allied [478]. allocation [71, 315]. alphabet [155]. alternating [383]. alternative [221]. among [3]. amplitude [454]. Analysis [126, 146, 187, 217, 218, 237, 262, 268, 297, 309, 348, 385, 444, 454, 457]. analytical [204, 465]. Analyze [198]. analyzing [185]. Angularly [112]. annihilators [336]. anonymous [168, 235, 397]. ant [279]. antenna [204]. Anti [64]. Anti- [64]. antimagic [403]. AODV [269]. Application [266, 329]. Applications [106, 297, 324, 350, 418, 441, 479]. applying [306]. approach [47, 53, 160, 202, 204, 251, 414, 423, 436, 437, 464, 473]. approaches [314, 365]. approximate [29, 33, 295]. approximation [414]. April [446]. arbitrary [94]. arcs [419]. areas [58]. arising [346]. arithmetic [399]. arithmetical [11]. Arnold [463]. arrangements [327]. arrays [73, 124, 167, 366]. ary [132, 155, 172, 175, 177, 454]. Ascents [298]. aspects [23]. assessment [338]. assessments [25]. assignment [111]. associated [94, 95, 208]. association [373]. assuming [62]. asymmetric [73]. Asymptotic [217, 218, 365, 410]. ATM [248]. attached [96]. attack [133, 194, 215, 364]. Attacks [146, 209, 227]. attribute [338]. attributes [459]. auction [235]. August [413, 460]. authenticated [239, 247]. Authentication [51, 57, 60, 63, 157, 192–194, 220, 364, 397]. authority [396]. automata [458]. automatic [381, 409]. automation [96]. automobile [251]. average [471]. aware [285]. AWGN [280].

backward [428]. Baer [101]. balanced [75, 80, 85, 154, 172, 177, 212, 283]. ballot [175]. ballot-sequences [175]. Banach [307]. based [27, 57, 69, 71, 72, 78, 79, 107, 110, 123, 158, 186, 192– 194, 205, 209, 215, 220, 236, 254, 265, 269, 312, 314, 341, 343, 353, 376, 381, 388, 390, 393, 396, 398, 414, 415, 421, 436, 439, 442, 447, 453, 457, 459, 461, 465]. based-model [393]. bases [166, 208]. basis [393]. beam [322]. behavior [381]. Benes [228]. Bent [184, 345]. Berge [362, 363]. between [41, 117, 270, 290, 359, 472]. bilinear [421]. binary [55, 282, 301, 395, 415]. binomial [217, 218, 418]. Biometric [423, 447]. bipartite [169, 340]. bit [188, 297, 407, 420]. biting [180, 387]. Black [43]. blict [52, 149]. Blind [51, 53, 106, 123]. blitact [52, 149]. Block [19, 47, 80, 127, 180, 185, 212, 288, 405, 406]. block-stepwise [185]. Blocking [1, 8, 61, 104, 252, 327]. bondage [5, 349]. Book [412]. boolean [75, 118, 187, 210, 211, 273, 274, 308, 345, 426, 439, 450]. Bound [45, 138]. bounds [118, 180]. box [136, 216]. boxes [255]. branching [62]. broadcast [189]. Bruijn [65]. Brute [227]. buckets [125]. Burkard [176]. burst [102, 138].

CAAA [459]. Calculating [242]. calculation [267]. calculus [348, 382]. Call [341, 371]. cameras [320]. Canonical [100]. Cao [364]. caps [380, 419]. case [45, 59]. cash [207]. categoric [41]. categories [88]. caterpillars [156]. Cayley [229, 274]. CDMA [254, 280]. cellular [315, 322, 341, 371, 458]. center [186]. centipedes [351]. centralizers [114]. certain [1, 166, 455]. changes [353]. channel [315, 385]. Characterization [2, 9, 16, 84, 101, 139, 170, 195, 372]. Characterizations [130, 257, 294, 380, 419]. Chinese [89]. chords [340]. chromatic [132, 156, 295]. cipher [110, 127, 271, 277, 405]. ciphered [407]. Ciphers [406, 458]. circuits [352, 435]. circulant [140]. circular [204]. class [13, 98, 143, 234, 291, 329, 335, 455, 469]. Classes [11, 270, 401]. classification [126, 151, 173, 405]. Classifications [76]. classifiers [408]. cliques [213]. closed [48]. closing [373]. clustering [442, 443, 448, 457, 461, 474]. clusters [457]. co [219, 367]. co-induction [219]. co-operating [367]. COA [122]. code [63, 108, 299, 311, 464, 468]. codes [55, 80, 82, 114, 144, 150, 154, 155, 172, 180, 182, 208, 214, 230, 231, 259, 333, 339, 387, 409, 440, 473]. Codeword [190]. coding [51, 154, 199, 410]. collective [76]. colony [279]. coloring [57]. Combinatorial [84, 101, 135, 139, 154, 270, 283, 372, 380, 441]. combined [55]. Comments [194]. Common [134, 245]. Communication [459]. Commutative [94, 134]. Comparative [126, 408]. Comparing [314]. comparison [322]. compiler [219]. complement [92, 327]. Complementary [141, 232, 293, 361]. complete [132, 156, 197, 340, 350, 435, 468]. completely [54]. completeness [355]. complex [154]. complexity [21, 177, 180, 282]. component [373]. composite [316]. Composition [265, 424]. compression [441]. computation [107, 117]. computational [126]. computationally [79]. computations [72]. computer [275, 338]. computerized [251]. Computing [24]. concatenated [180]. concepts [399]. concurrent [107]. condition [176, 475]. conditional [35]. conditions [38]. cones [9]. conference [283]. configuration [16]. configurations [50]. conflict [309]. congruential [188, 420]. conjecture [296, 300, 316, 362, 363]. Connected [171]. Connections [41, 359]. connectivity [400]. consistent [320]. consonant [128]. Construct [63]. constructed [214]. constructing [72, 243, 283]. construction [73, 90, 136, 140, 177, 212, 284, 391, 401, 409]. Constructions [135, 350]. containing [83]. Contents [378, 394, 404, 413, 422, 431, 438, 446, 452, 460, 467, 477]. Control [22, 36, 78, 111, 196, 341]. controllability [108]. convergence [237]. Convolution [384]. cooperative [36]. core [279]. cores [140]. correcting [409]. correction [55, 138]. correctness [219]. correlation [146, 284]. corruption [31]. cosets [64]. cototal [18]. counting [234]. cover [115, 162]. covering [68, 82, 118, 264, 347, 350, 449]. Coverings [435]. criterion [429]. critical [344, 427, 462]. crossing [10, 197]. Cryptanalysis [439]. crypto [236]. cryptographic [111, 124, 216, 398, 434]. Cryptography [121, 158, 247, 263, 286, 289, 292, 312]. cryptology [153]. cryptosystem [205, 215, 379, 408, 439]. Cryptosystems [66, 103, 375]. Cubic [142, 257, 294, 414]. cubics [116, 179]. cubs [83]. curvature [414]. curve [153, 247, 250, 414, 416]. curve-shaped [414]. curves [142, 257, 396, 453]. customer [341]. cybernetics [346]. cycle [340, 357]. cyclic [333, 339]. cyclically [455].

D [160]. data [70, 77, 112, 203, 266, 341, 415, 447]. December [431, 477]. decision [32, 338]. decoding [180, 190]. decomposable [54]. Defense [381]. defining [163]. defuzzification [122]. Degree [174, 332, 462, 465]. Degrees [462]. delay [71]. dematerialization [423]. deniable [397]. densities [418]. density [102, 385]. derivable [80]. derivation [241]. derivations [6]. derivative [11]. derived [124, 469]. descents [298]. Design [78, 172, 212, 251, 252, 283, 299, 398]. designated [206]. designs [8, 80, 113, 135, 161, 166, 185, 409]. detecting [353]. Detection [55, 102, 160, 164, 183, 279, 353, 436, 437, 442, 443, 465]. determinant [3]. determinants [17]. determining [62]. Development [3, 309, 368]. diagnosis [32]. Diameter [105, 249]. difference [384, 392]. differences [3]. differential [469]. difficult [362]. Diffraction [70, 77, 112]. digital [69, 266]. Dimension [210, 211]. dioids [386]. direct [7, 393]. directed [421, 435]. Dirichlet [14]. Discrete [26, 59, 69, 121, 195, 220, 343, 346, 376, 414, 478]. dissertation [363]. distance [46, 259, 471]. distinguishability [190]. distortion [51]. Distributed [36, 410]. distribution [34, 324]. distributions [23]. distributive [130]. diversity [190]. divisible [161, 207]. divisive [474]. divisive-agglomerative [474]. DNA [72, 107, 289]. documents [27]. domains [221]. domestic [472]. Dominating [351, 357]. Domination [18, 174, 186, 232, 240, 244, 245, 326, 334, 356, 451, 470]. double [444]. Doubly [182, 426]. doubly-even [182]. DRC [347]. DS [254, 280]. DS-CDMA [254, 280]. DSA [303]. dual [86, 182, 214, 297, 468]. DWT [436]. Dynamic [181, 315, 437, 448, 461].

each [3]. ECM [329]. economies [34]. Edge [56, 65, 225, 226, 246, 449, 462]. educational [337]. EEG [126]. effective [117]. Efficient [154, 155, 167, 193, 235, 352, 397, 433, 474]. Eigenvalues [392, 400]. Electronic [168, 207, 396, 447]. elements [3, 242, 336]. elliptic [67, 153, 247, 250, 396, 416, 453]. embeddability [191]. Embedding [169, 331, 355]. Encryption [147, 203, 266, 288, 366, 376, 406]. engineering [423]. English [128]. enhanced [194]. Enhancement [30, 196, 220]. entries [358]. entropy [6, 62]. Enumeration [383]. Enumerations [197]. enumerator [468]. environments [122, 193, 280]. Ephemeral [382]. equations [131, 149, 402, 469]. equitable [317]. equivalence [290]. equivalent [389]. Errata [310]. error [55, 102, 138, 177, 180, 282, 409]. Essentially [60]. establishment [330]. estimation [341]. evaluating [122]. evaluation [280]. even [182, 427]. events [25]. Every [109, 427]. evolution [385, 469]. evolutionary [313]. Exact [29, 132]. excellent [109]. exceptions [111]. exchange [247, 391]. exhaustive [444, 456]. existence [1]. expansion [217, 218, 433]. experiments [409]. Exploring [114]. Exponential [180]. expression [160, 308]. Extended [181, 286, 303]. extending [163]. Extension [301]. extensions [34]. Extremal [37, 75, 182].

Face [160, 314, 390]. facial [160]. factorial [23]. factoring [69, 209, 302, 343, 376]. factorization [123]. factorizations [375]. Fail [123, 281]. Fail-stop [123, 281]. families [134]. Fano [83]. Farey [22]. Fast [146, 360]. Fault [215]. feasibility [47]. features [390]. February [438]. Fermat [224]. Fibonacci [42, 267, 270, 272]. field [159, 358]. fields [214]. fighter [306]. Film [472]. finding [222, 224]. Fingerprint [436]. finite [45, 120, 184, 242, 358, 385]. finite-state [385]. first [299, 379]. fixed [134, 307]. flexible [109, 168]. floating [444]. fold [60]. force [227]. forcing [449]. formulae [365]. fortification [93]. four [311]. Fourier [26, 187]. fractions [22]. frauds [51]. free [427]. friendly [233, 316]. function [3, 110, 297, 393, 450]. functional [131]. functions [11, 14, 75, 118, 157, 184, 187, 195, 210, 211, 217, 218, 227, 273, 274, 330, 345]. Fundamentals [199]. Further [268, 305, 445]. Fusion [447]. fuzziness [4]. Fuzzy [7, 25, 41, 58, 64, 122, 320, 359, 388, 395, 399].

Galois [9, 24]. general [207, 373]. Generalised [463]. generalization [17, 260]. generalizations [165]. Generalized [42, 50, 71, 148, 203, 328, 384, 387, 414, 468]. generated [455]. generating [217, 218]. Generation [175, 250, 308]. generator [411, 420]. generators [114, 188, 428, 444, 456]. genesis [17]. Genetic [140, 233, 313]. geodetic [225]. Geodomination [171, 321]. geometric [160]. geometry [126]. Germain [316]. Gilbert [45]. Goldbach [296, 316]. golden [159]. good [456]. governed [36]. graceful [291, 300, 355, 417]. Gracefulness [340]. Graph [5, 18, 44, 48, 52, 68, 86, 109, 141, 149, 169, 174, 225, 226, 229, 243, 244, 246, 274, 276, 290, 295, 328, 332, 334, 349, 370, 427, 429, 432, 449, 451, 471]. Graphic [9, 257, 294, 419]. graphical [17]. graphoidal [28, 68, 115]. Graphs [10, 12, 19, 28, 41, 46, 56, 65, 105, 115, 149, 151, 156, 163, 171, 176, 197, 213, 232, 240, 249, 256, 261, 278, 300, 301, 305, 317, 321, 323, 326, 340, 342, 344, 354– 356, 359, 374, 375, 400, 401, 417, 425, 435, 445, 462]. gray [333]. grey [262]. Gröbner [166, 208]. group [161, 184, 250]. groups [242, 359, 455]. GSIC [254]. guessing [364].

H3 [455]. H4 [455]. Hadamard [125, 140, 182, 191, 214]. Hadwiger [363]. Hamiltonian [107, 229, 261, 400, 475]. Hammer [176]. Hanoi [241]. Hard [416]. hardware [353]. hash [110, 157, 227, 297]. hashing [128]. heuristic [428]. Hidden [318]. hiding [79, 221, 415]. hierarchical [196, 474]. hierarchy [111, 189]. High [102, 155]. High-density-burst [102]. high-order [155]. highly [30]. Hill [277]. Histogram [406]. historical [17]. hoc [289]. honeycomb [470]. hubs [443, 448, 465, 466]. HV [293]. HV-rings [293]. hyper [87, 173, 183]. hyper-Latin [183]. hypergraphs [41, 85]. Hypergroupoids [66, 91]. hypergroups [88, 90, 92–94, 96, 100, 104, 395]. hyperideals [200, 201, 388]. hypermodules [88]. hyperoperations [95]. hyperplane [327]. hyperrings [97, 98]. hyperstructures [89, 293, 479]. hypothesis [318].

ICA [254]. ID [193, 421]. ID-based [193, 421]. IDEA [110]. ideals [87]. identically [410]. Identification [126, 167, 193, 405–408]. identifying [278]. Identikit [265]. identities [93, 195, 434]. identity [312, 469]. identity-based [312]. II [77, 321]. III [112]. Image [27, 202, 313, 333, 415, 441, 461]. imaging [70, 77, 112]. Implication [388]. implicative [87]. Improved [15, 69, 267, 279, 313]. improvement [146]. improving [379]. improvised [277]. incidence [20]. incomplete [80, 212]. independent [410]. Index [78, 258, 319, 348, 377, 382, 430, 476]. Indian [407]. Indices [216]. induction [219]. industry [338]. Inequivalent [125]. inferences [76]. informatics [346]. Information [53, 79, 131, 221, 288, 318, 337, 447]. infosecurity [423]. inner [180]. innovative [447]. insecure [239]. integer [123, 223]. integers [223, 324]. Integral [275, 337]. integrals [35]. integrated [341, 437]. Interaction [183]. interactive [205, 397]. Interchange [354]. interconnected [393]. interior [388]. interleaved [284]. intersection [370, 425]. Intersections [113]. interval [359]. intractability [57]. intractable [79]. intrusion [353]. invariant [114, 172]. Inverse [240]. inverses [129, 325]. inversion [70, 77]. involutary [450]. involution [230, 231, 402, 440]. involutions [229]. involving [375, 447]. IP [248]. IPv6 [153]. irreducible [16, 83, 116, 179]. isomorphic [167]. isomorphism [270]. isomorphisms [117]. iterative [15, 146].

J [121]. JERIM [297]. JERIM-320 [297]. joined [323]. joint [282]. judgements [49]. June [404, 452].

Kernel [108]. Key [22, 111, 133, 189, 203, 220, 236, 239, 247, 285, 292, 330, 364, 376, 382, 391, 439, 447, 469]. Key-lock [22]. Keyed [81, 157]. Keyed/unkeyed [81]. keying [103]. keys [128, 307, 329]. kind [370]. kinds [87]. King [204]. Klein [59]. Krawtchouk [468].

labeling [137, 256, 276, 278, 317, 320, 403]. Laceability [46]. Landau [157]. Landmark [265]. Landmark-based [265]. Landmarks [178]. lane [264]. language [402]. languages [29, 150, 407]. Laplacian [326, 400]. large [1, 177]. last [362]. Latin [183, 398]. Lattice [386]. lattices [222, 336]. layout [228]. LDA [314]. leakage [288]. Lee [138, 392]. left [434]. length [383]. less [34]. lesser [238, 361]. letter [128]. letter-oriented [128]. level [60, 135, 166]. Li [196]. LICoS [265]. Lict [226]. light [166]. limitation [145]. limited [112]. line [3, 149, 255, 301, 364]. line-splitting [301]. linear [21, 40, 47, 114, 177, 188, 208, 242, 282, 420, 473]. lines [61, 139]. Linux [127]. local [437, 448]. locating [169]. location [285]. location-aware [285]. lock [22]. logarithm [376]. logarithms [69, 220, 343]. logic [320]. long [47, 210, 211]. loop [204]. loops [434]. lossless [410, 441]. Low [284, 288]. LSB [169]. Lu [364]. Lucas [272].

machine [368]. magic [137]. Major [472]. making [338]. malicious [381]. management [189, 285, 312]. MANETs [289]. Markov [385]. massively [72]. matching [29, 107, 357, 427]. Math [121]. Mathematical [241, 309]. mathematics [478]. matrices [17, 125, 129, 140, 182, 214, 325, 358, 392]. matrix [29, 203, 268, 274, 360]. matroids [301]. maximal [176]. maximally [75]. Maximum [83, 174, 197, 331]. Mean [256, 278, 401, 445]. measurable [131]. measure [216]. measures [62]. mechanism [22, 196]. medical [32]. Mersenne [316]. message [63, 157, 198, 318]. Meta [457]. metacirculant [151]. Method [15, 122, 136, 185, 262, 348, 382, 406, 415, 461]. methodology [152]. methods [126, 212]. micro [448]. Minimal [1, 8, 72, 274]. minimally [19]. minimization [78]. Minimum [115, 190, 437, 442, 443, 448, 457, 461, 465, 466, 474]. mining [461]. mixed [124, 135, 371]. mixed-level [135]. Mixer [271]. mobile [193, 289, 315, 322, 367, 371]. modal [464]. mode [128, 147]. model [31, 34, 59, 71, 318, 393]. models [70, 275, 290, 337]. modified [23, 393]. modulation [454]. moduli [302]. modulus [411, 420]. morphic [402]. morphisms [90]. Motion [320]. Muller [299]. Multi [107, 197, 235, 241, 312]. Multi-identity [312]. multi-partite [197]. multi-peg [241]. multi-round [235]. multi-string [107]. multi-text [107]. multigraph [350]. Multiple [307, 320, 338, 411, 428, 444, 456]. multiplication [360, 416]. Multiplicative [11, 14, 97, 293, 336]. multipliers [188, 444]. multirate [254]. multisequences [282]. mutants [308]. mutual [193, 239].

navigation [381]. necessary [176]. neighbourhood [48]. Nested [80, 161]. network [126, 248, 322, 341, 353, 373, 393]. Networks [178, 228, 239, 285, 289, 315, 371, 470]. Neural [126, 353, 373, 393]. Nimix [450]. no [121]. noisy [30]. non [1, 8, 205, 304, 397, 429]. non- [429]. non-existence [1]. non-interactive [205, 397]. non-symmetric [8]. nonabelian [184]. nonhamiltonian [176]. nonlinear [34, 75, 255, 426, 450]. nonouterplanar [19]. Nonparametric [466]. normal [64]. norms [359]. Note [2, 42, 68, 85, 91, 141, 311, 334, 344, 387, 402, 409, 432, 471]. notebook [338]. notes [272]. nothing [34]. notion [464]. novel [269, 289, 369, 461]. NP [269, 355]. NP-based [269]. NP-completeness [355]. null [155]. Number [18, 44, 68, 83, 132, 156, 188, 222, 225, 244–246, 249, 295, 310, 328, 334, 349, 356, 374, 378, 394, 404, 411, 418, 420, 422, 428, 431, 444, 449, 456]. numbers [5, 10, 42, 58, 197, 267, 316, 326, 365, 447, 463, 465]. numeral [415]. Numerical [11, 204].

objects [320, 414]. oblivious [205, 221]. observer [78]. observer-based [78]. obtained [278, 395]. October [422, 467]. odd [59, 383, 435]. On-line [364]. one [106, 110, 330, 375]. one-time [106]. one-way [110, 330]. online [286]. operating [127, 367]. operation [153, 268, 301]. operations [275]. operator [36]. optimal [135, 169, 243, 350]. optimality [38]. optimization [219, 279]. optimized [393]. order [38, 59, 96, 155, 173, 299, 311, 411, 427, 433, 456]. ordered [90, 261]. orders [242]. oriented [128]. original [316]. Orthogonal [39, 73, 124, 167, 287, 366, 409]. Osborn [434]. outlier [437, 442]. outliers [443, 448, 465, 466].

packet [71]. packing [350]. packings [435]. padding [288]. pair [22]. pairings [421]. pairs [224]. parabolic [372]. paradigm [289, 367]. paradox [34]. Parallel [72, 175, 243, 283, 306, 340]. parallelism [113]. parameter [12]. parameters [174, 276]. part [275, 337, 362]. partial [367, 456]. partite [197]. partition [392]. partitioning [202, 223]. partitions [324]. password [194, 364]. Path [105, 107, 162, 369]. paths [229, 340]. patterns [234]. pay [189]. pay-TV [189]. PDE [433]. Pebbling [40, 328]. peg [241]. pentadiagonal [325]. Perfect [255, 259, 316, 357, 426, 427]. Performance [127, 248, 280, 322, 338, 341, 348, 387, 408, 454, 457]. Performing [373]. Periphery [213]. permanent [342]. permutation [172, 234, 300]. permutations [17, 43, 148, 272, 273, 383, 426, 439]. person [126]. perspective [465]. piecewise [410]. plain [407]. planar [226]. Planarity [323, 429]. plane [83, 120, 306, 311]. planes [2]. point [51, 162, 245, 279, 307, 444]. points [83, 134]. Polya [324]. polygroups [64]. polynomial [274]. Polynomials [37, 97, 152, 424, 433, 468]. poset [130]. positive [87, 223]. possessing [62]. Power [132, 156, 170, 470]. precedents [17]. precision [444]. prediction [373]. preferences [76]. prefix [164, 170]. presented [455]. prevention [31]. price [373]. primary [7]. prime [7, 214, 332, 365, 420, 447]. primes [159, 238, 316, 361]. Primitive [152]. principle [373, 424]. privacy [389]. Private [244]. probabilistic [365]. probability [4, 25, 55]. problem [33, 47, 57, 107, 151, 186, 264, 363]. problems [37, 79, 199, 324, 416]. procedure [49]. processes [337, 346]. processing [26, 27]. product [21]. profile [353]. programming [40]. projections [47]. projective [2, 311, 327]. properties [119, 141, 154, 234, 250, 386, 440]. property [62, 410]. protected [209]. protection [367, 397]. protocol [57, 235, 247, 269, 330, 364, 391, 397, 469]. protocols [459]. Provably [209, 330]. Proxy [145, 209, 281, 421]. proxy-protected [209]. pseudorandom [250]. Public [236, 292, 376, 423, 439]. public-key [236]. purposes [126].

quadratic [78, 159]. quadrature [454]. quadric [9, 84]. quadrics [67, 372]. quantum [352]. quartics [304]. Quasi [10, 91, 395]. quasi-hypergroups [395]. quasi-Steiner [91]. Quasi-total [10]. quasigroups [61]. quaternion [26].

Rabin [302]. radial [393]. radicals [20]. radii [118]. Radio [249, 374]. radius [82]. rake [280]. Ramanujan [157]. Random [77, 188, 275, 288, 290, 298, 346, 411, 420, 428, 444, 456]. randomized [203]. Randomizing [147]. randomly [360]. rank [259]. Ranking [58, 390]. rate [51, 71]. rate-based [71]. rate-distortion [51]. ratio [49, 58]. Re [103]. Re-keying [103]. real [255]. real-line [255]. realistic [207]. receiver [254, 280]. reciprocal [24]. recognition [160, 314, 390]. reconstruction [56]. recovery [382]. rectilinear [197]. recurring [21]. recursive [411, 428, 444, 456]. Rédei [61, 165]. Reed [299]. refinement [15]. reformulation [296]. region [390]. region-based [390]. regression [373, 405]. Regular [44, 86, 163]. reinvestigated [389]. related [230, 231, 358]. Relation [472]. relations [395]. relative [58, 336]. releasing [472]. remark [253]. rentals [472]. Renyi [260]. Representation [62, 159, 352]. residue [236]. resolution [77, 309]. respect [213]. Restrained [349]. restricted [444]. result [260]. results [43, 305, 361, 367, 445]. retrials [341]. Retrieval [265]. reverse [423]. Reviews [412]. revocation [281]. Rijndael [215]. ring [14]. rings [24, 236, 293]. Robust [205]. romanized [407]. rooted [291]. roots [154]. round [235]. routes [222]. routing [269, 289]. row [360]. RSA [205, 302, 329]. rule [17]. runs [210, 211].

S [136, 216, 255]. S-box [136, 216]. S-boxes [255]. Sarrus [17]. scalar [416]. scale [34]. schedule [198, 283]. scheduling [71]. scheme [69, 110, 111, 123, 158, 168, 193, 194, 203, 221, 234, 239, 269, 281, 285, 286, 288, 343, 371, 376, 407, 421]. schemes [17, 124, 145, 209, 220, 341, 366, 398]. Schrodinger [36]. Sci [121]. scope [34]. SDMA [322]. search [238, 361, 456, 466]. Second [38, 433]. secret [169, 206]. section [159]. Secure [60, 79, 144, 209, 285, 289, 330, 364]. secured [313]. securing [27]. Security [196, 220, 269]. Segrè [253]. selected [360]. selective [415]. Self [182, 214, 354]. Self-dual [182, 214]. semientire [226]. semigroup [292]. semihypergroups [388]. semihyperrings [200, 201]. Semiprime [336]. semirings [292]. semitotal [19]. semitotal-block [19]. sender [397]. sensor [285]. separation [53]. sequence [177, 250, 462]. sequences [21, 175, 284]. sequential [355]. series [23, 217, 218, 463]. session [283]. set [242, 252, 305, 355]. set-graceful [355]. set-sequential [355]. set-valued [305]. sets [1, 8, 13, 61, 74, 104, 120, 121, 143, 195, 230, 231, 284, 327, 335, 351, 384]. SFP [144]. SHA [63, 81, 198]. SHA-1 [198]. SHA-2 [81]. Shannon [6]. shaped [414]. sharing [206]. Sharpening [45]. shift [284]. shortest [222, 369]. shuffling [202]. SIFT [314, 390]. signal [26, 284]. signals [30]. signature [69, 106, 123, 145, 206, 209, 281, 287, 343, 421]. Signed [326, 425]. similar [354]. similarity [443, 448]. simple [243, 363]. simplex [80, 82]. simplification [362]. simultaneous [47]. Single [369, 396]. singular [116, 179, 304]. singularity [217, 218]. size [252]. sizes [54]. skew [139, 214]. skew-Hadamard [214]. solid [138]. solution [33, 433, 469]. solutions [131, 402]. Some [6, 23, 34, 37, 41, 43, 73, 74, 87, 116, 121, 127, 131, 135, 165, 214, 234, 257, 272, 302, 356, 358, 365, 399, 418, 440]. Sophie [316]. source [53, 369]. sources [410]. space [4, 9, 327, 370]. spanning [72, 243, 437, 442, 443, 448, 457, 461, 465, 466, 474]. spatial [221]. special [234]. specific [185]. spectral [155]. spectral-null [155]. spectrum [161, 163, 326]. speech [30]. Split [5, 334]. splitting [301]. square [17, 164, 222, 229]. squares [159, 183, 398]. state [385]. static [285]. statistical [297]. Steiner [91, 120, 246]. steps [47]. stepwise [185]. stock [373]. stop [123, 281]. strategy [472]. stream [271, 407, 458]. strength [216]. strengthening [458]. string [107, 164]. Strong [93, 163, 398]. Structural [443, 448]. structure [78, 208]. structures [90, 95, 99, 158, 286]. STS [158, 286]. students [457]. study [127, 204, 212, 261, 306, 379]. sub [64]. sub- [64]. subgroups [41]. subhypergroups [41]. subhypermodules [7]. sublinear [164]. subsets [41]. subspaces [101]. substitution [133, 169, 424]. sufficient [475]. sum [7, 223]. sums [65, 159, 275, 346]. super [77, 137, 401, 445]. super-resolution [77]. supersaturated [135, 185]. Support [32, 54, 405]. survey [56, 347]. Suspect [265]. suspension [251]. switch [368]. switchings [354]. symbol [172]. symmetric [8, 73, 103, 127, 252, 307]. symmetry [62]. synthesizing [49]. system [32, 127, 236, 251, 306, 353, 415]. Systematic [202]. systemics [346]. systems [26, 36, 54, 60, 120, 189, 254, 275, 337].

Table [394]. tact [368]. tactile [368]. tail [180, 387]. tank [393]. Taylor [433]. TCP [248]. TCP/IP [248]. technique [203, 206, 447]. techniques [27]. term [428, 444, 456]. testing [152, 318, 368]. tetravalent [151]. text [107]. th [154]. THBL [306]. their [25, 90, 163, 250, 407]. theorem [35, 165, 253, 307, 324]. theorems [355]. theoretic [53, 318]. Theoretical [118]. Theory [51, 131, 172, 418, 473]. third [456]. third-order [456]. three [249, 283, 322, 393, 434]. three-parallel [283]. threshold [119, 221, 236]. time [106, 145, 164, 264]. timestamp [194]. timestamp-based [194]. tomographic [70, 77, 112]. tools [381]. topics [478]. topologically [355]. torus [178]. total [10, 19, 48, 232, 240, 326, 432]. total-block [19]. Tower [241]. traceable [475]. tracking [320]. traffic [322, 371]. transfer [205]. transferable [207]. transform [26]. transformation [263, 429, 451]. transformations [134, 192]. transforms [441]. Translation [35]. transportation [33]. Transposition [92, 93, 379]. trapezoid [243]. tree [132, 141, 243, 437, 442, 443, 448, 457, 461, 465, 466, 474]. trees [72, 175, 291, 298]. trellis [180, 387]. Triangular [256, 454]. tribes [22]. tridiagonal [129]. triple [54, 120]. TV [189]. twin [316]. twisted [294]. two [61, 139, 140, 159, 166, 278, 322, 428, 444, 456]. two-term [428, 444, 456]. type [74, 120, 121, 161, 185, 199].

U.S. [472]. unbalanced [33]. unconditionally [285]. undetected [55]. Uniform [13]. union [340]. unique [115]. units [24]. unity [154]. universal [410, 434]. unkeyed [81]. upon [221]. Upper [118, 279]. usage [458]. use [207, 373]. used [55]. useful [131]. User [57, 193, 233]. User-friendly [233]. Using [22, 40, 166, 203, 206, 217–219, 262, 279, 292, 307, 313, 320, 330, 366, 367, 382, 390, 405, 406, 409, 415, 433, 437, 443, 448, 465, 466, 474].

valent [464]. valued [305, 359]. values [132]. variables [273]. Various [17]. Varshamov [45]. VC [211]. VC-dimension [211]. vector [360, 370, 405]. vectorial [450]. Verification [219]. verifier [206]. Veronesean [84]. version [277]. Vertex [137, 317, 344, 354, 417, 427]. Vertex-graceful [417]. vertex-magic [137]. via [126, 318, 423]. view [51]. vision [320]. visual [124, 158, 286]. VLSI [228]. voice [341]. voice/data [341]. Vol. [438, 446, 452, 460, 467, 477]. Volume [258, 310, 319, 377, 378, 394, 404, 413, 422, 430, 431, 476]. voting [168, 396]. Vowel [128]. Vowel-consonant [128].

Walsh [187]. water [393]. watermarking [221, 423]. wavelet [192, 423]. wavelets [39]. way [110, 330]. weakly [7]. weight [138, 415, 468]. weight-based [415]. Weighted [169]. Weighting [262]. Which [104, 172]. Wiener [35]. windows [264]. Wired [389]. wireless [248, 341]. without [62]. WLT [194]. words [170].

Yang [196]. Yeh [35].

References

Innamorati:1998:NEC

[1] Stefano Innamorati. The non-existence of certain large minimal blocking sets. Journal of Discrete Mathematical Sciences and Cryptography, 1(1): 1–26, 1998. CODEN ???? ISSN 0972-0529.

Zannetti:1998:NCP

[2] Mauro Zannetti. Note on a characterization of projective planes. Journal of Discrete Mathematical Sciences and Cryptography, 1(1):27–28, 1998. CODEN ???? ISSN 0972-0529.

Monaco:1998:DDF

[3] Giuseppe Monaco and Antonio Monaco. Development of a determinant in function of the differences among the elements of each line. Journal of Discrete Mathematical Sciences and Cryptography, 1(1):29–47, 1998. CODEN ???? ISSN 0972-0529.

Prasad:1998:FPS

[4] S. K. Prasad. On fuzziness in a probability space. Journal of Discrete Mathematical Sciences and Cryptography, 1(1):73–77, 1998. CODEN ???? ISSN 0972-0529.

Kulli:1998:SBN

[5] V. R. Kulli, B. Janakiram, and Radha R. Iyer. Split bondage numbers of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 1 (1):79–84, 1998. CODEN ???? ISSN 0972-0529.

Nath:1998:SDS

[6] Prem Nath, M. M. Kaur, and Ranjeet Singh. Some derivations of the Shannon entropy. Journal of Discrete Mathematical Sciences and Cryptography, 1(1):85–100, 1998. CODEN ???? ISSN 0972-0529.

Ameri:1998:FDS

[7] R. Ameri and M. M. Zahedi. Fuzzy direct sum and fuzzy (weakly) prime and primary subhypermodules. Journal of Discrete Mathematical Sciences and Cryptography, 1(2-3):101–116, 1998. CODEN ???? ISSN 0972-0529.

Berardi:1998:MBS

[8] Luigia Berardi and Fulvio Zuanni. On minimal blocking sets in nonsymmetric designs. Journal of Discrete Mathematical Sciences and Cryptography, 1(2-3):117–125, 1998. CODEN ???? ISSN 0972-0529.

Ferri:1998:GCQ

[9] Osvaldo Ferri and Stefania Ferri. Graphic characterization of quadric cones in a Galois space PG(3, q). Journal of Discrete Mathematical Sciences and Cryptography, 1(2-3):127–131, 1998. CODEN ???? ISSN 0972-0529.

Basavanagoud:1998:QTG

[10] B. Basavanagoud. Quasi-total graphs with crossing numbers. Journal of Discrete Mathematical Sciences and Cryptography, 1(2-3):133–142, 1998. CODEN ???? ISSN 0972-0529.

Palumbo:1999:CAM

[11] Biagio Palumbo. Classes of arithmetical multiplicative functions and numerical derivative. Journal of Discrete Mathematical Sciences and Cryptography, 2(1):1–20, 1999. CODEN ???? ISSN 0972-0529.

Gionfriddo:1999:PG

[12] Lucia Gionfriddo. On the parameter m_{2}(h) for graphs. Journal of Discrete Mathematical Sciences and Cryptography, 2(1):31–34, 1999. CODEN ???? ISSN 0972-0529.

Abatangelo:1999:USC

[13] L. Maria Abatangelo and Manuela Erroi. Uniform 20-sets of class [0, 2, 3, 4, 5] in AG(2, 7). Journal of Discrete Mathematical Sciences and Cryptography, 2(1):35–44, 1999. CODEN ???? ISSN 0972-0529.

diGennaro:1999:MFD

[14] Fernando di Gennaro and Salvatore Furneri. On C_{λ} and F_{λ} multiplicative functions in Dirichlet ring. Journal of Discrete Mathematical Sciences and Cryptography, 2(1):59–72, 1999. CODEN ???? ISSN 0972-0529.

Al-Towaiq:1999:IAI

[15] M. H. Al-Towaiq. An improved algorithm for the iterative refinement method. Journal of Discrete Mathematical Sciences and Cryptography, 2 (1):73–76, 1999. CODEN ???? ISSN 0972-0529.

Teodori:1999:CIC

[16] Nicola Teodori. A characterization of irreducible configuration (n_{3}). Journal of Discrete Mathematical Sciences and Cryptography, 2(1):77–83, 1999. CODEN ???? ISSN 0972-0529.

Monaco:1999:VGP

[17] Giuseppe Monaco and Antonio Monaco. Various genesis of permutations; determinants; square matrices as graphical schemes; generalization of Sarrus’ rule and its historical precedents. Journal of Discrete Mathematical Sciences and Cryptography, 2(2-3):141–177, 1999. CODEN ???? ISSN 0972-0529.

Kulli:1999:CDN

[18] V. R. Kulli, B. Janakiram, and Radha R. Iyer. The cototal domination number of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 2(2-3):179–184, 1999. CODEN ???? ISSN 0972-0529.

Akka:1999:MNS

[19] D. G. Akka and M. S. Patil. 2-minimally nonouterplanar semitotal-block graphs and total-block graphs. Journal of Discrete Mathematical Sciences and Cryptography, 2(2-3):185–196, 1999. CODEN ???? ISSN 0972-0529.

Spiegel:1999:ARI

[20] Eugene Spiegel. Algebraic radicals and incidence algebras. Journal of Discrete Mathematical Sciences and Cryptography, 2(2-3):197–204, 1999. CODEN ???? ISSN 0972-0529.

Morgari:2000:LCP

[21] Guglielmo Morgari, Oscar Steila, and Michele Elia. On the linear complexity of product sequences of linear recurring sequences. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):1–9, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Chandrashekhar:2000:KLP

[22] H. Chandrashekhar and M. Nagaraj. Key-lock pair mechanism for access control using tribes of Farey fractions. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):11–22, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Das:2000:SAM

[23] Kishore K. Das and Nibedita Bhattacharjee. Some aspects of modified factorial series distributions. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):23–40, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Elia:2000:CRU

[24] Michele Elia, J. Carmelo Interlando, and Reginaldo Palazzo, Jr. Computing the reciprocal of units in Galois rings. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):41–55, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Maturo:2000:FET

[25] Antonio Maturo. Fuzzy events and their probability assessments. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):83–94, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Chandrashekhar:2000:DQF

[26] H. Chandrashekhar and M. Nagaraj. Discrete quaternion Fourier transform in signal processing systems. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):95–112, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Ratan:2000:IPB

[27] Ram Ratan and P. K. Saxena. Image processing based techniques for securing documents. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):113–129, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Arumugam:2000:GG

[28] S. Arumugam and J. Suresh Suseela. On graphoidal graphs. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):131–137, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Radhakrishnan:2000:EAM

[29] V. Radhakrishnan, T. C. Venkatesan, and K. Krithivasan. Exact and approximate matching of matrix languages. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):139–156, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Pal:2000:EHN

[30] S. K. Pal and P. K. Saxena. Enhancement of highly noisy speech signals. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):157– 172, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Khan:2000:CPM

[31] Md. Asim Uddin Khan. The corruption prevention model. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):173–178, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Mukhopadhyay:2000:DSS

[32] S. Mukhopadhyay. A decision support system for medical diagnosis. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):179–192, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Mukhopadhyay:2000:ASU

[33] S. Mukhopadhyay and M. K. Singh. An approximate solution of an unbalanced transportation problem. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):193–202, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Ryan:2000:DMM

[34] Michael J. Ryan. The distribution model the more for less (nothing) paradox and economies of scale and scope: some nonlinear extensions. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):203– 211, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Chang:2000:TTC

[35] Kun Soo Chang, Sunkyung Kim, and Il Yoo. Translation theorem of conditional Yeh–Wiener integrals. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):213–225, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Serag:2000:DCC

[36] H. M. Serag. Distributed control for cooperative systems governed by Schrodinger operator. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):227–234, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Bidkham:2000:SEP

[37] M. Bidkham and K. K. Dewan. Some extremal problems of polynomials. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):235– 242, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Cambini:2000:SOO

[38] A. Cambini and L. Martein. Second order optimality conditions. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):243–252, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Mittal:2000:OW

[39] R. C. Mittal. Orthogonal wavelets. Journal of Discrete Mathematical Sciences and Cryptography, 3(1-3):253–262, 2000. CODEN ???? ISSN 0972-0529. Discrete mathematics and allied topics.

Lourdusamy:2001:PUL

[40] A. Lourdusamy and S. Somasundaram. Pebbling C_{5} ×C_{5} using linear programming. Journal of Discrete Mathematical Sciences and Cryptography, 4(1):1–15, 2001. CODEN ???? ISSN 0972-0529.

Zahedi:2001:SCC

[41] M. M. Zahedi and M. R. Khorashadi-Zadeh. Some categoric connections between fuzzy hypergraphs, subhypergroups, graphs, subgroups and subsets. Journal of Discrete Mathematical Sciences and Cryptography, 4(1): 17–32, 2001. CODEN ???? ISSN 0972-0529.

Eugeni:2001:NGF

[42] Franco Eugeni and Raffaele Mascella. A note on generalized Fibonacci numbers. Journal of Discrete Mathematical Sciences and Cryptography, 4 (1):33–45, 2001. CODEN ???? ISSN 0972-0529.

Foundas:2001:SRB

[43] E. Foundas. Some results of Black’s permutations. Journal of Discrete Mathematical Sciences and Cryptography, 4(1):47–55, 2001. CODEN ???? ISSN 0972-0529.

Kulli:2001:RNG

[44] V. R. Kulli, B. Janakiram, and Radha. R. Iyer. Regular number of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 4 (1):57–64, 2001. CODEN ???? ISSN 0972-0529.

Fabris:2001:SGV

[45] Francesco Fabris. Sharpening the Gilbert–Varshamov bound in the finite case. Journal of Discrete Mathematical Sciences and Cryptography, 4(1): 65–75, 2001. CODEN ???? ISSN 0972-0529.

Murali:2001:LDG

[46] R. Murali and K. S. Harinath. Laceability and distance graphs. Journal of Discrete Mathematical Sciences and Cryptography, 4(1):77–86, 2001. CODEN ???? ISSN 0972-0529.

Ozaktas:2001:ALS

[47] Hakan Özaktaş, Mustafa Ç. Pınar, Mustafa Akgül, Tahsin Kurç, and Cevdet Aykanat. An algorithm with long steps for the simultaneous block projections approach for the linear feasibility problem. Journal of Discrete Mathematical Sciences and Cryptography, 4(1):87–108, 2001. CO-DEN ???? ISSN 0972-0529.

Kulli:2001:TCN

[48] V. R. Kulli and Nanda S. Warad. On the total closed neighbourhood graph of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 4(2-3):109–114, 2001. CODEN ???? ISSN 0972-0529.

Hsueh:2001:PSR

[49] Kuang hua Hsueh, Patrick S. Chen, Peter Chu, and John Kwan han Liu. The procedure for synthesizing ratio judgements. Journal of Discrete Mathematical Sciences and Cryptography, 4(2-3):115–120, 2001. CODEN ???? ISSN 0972-0529.

Eugeni:2001:GC

[50] Franco Eugeni and Mauro Zannetti. Generalized configurations. Journal of Discrete Mathematical Sciences and Cryptography, 4(2-3):121–131, 2001. CODEN ???? ISSN 0972-0529.

Sgarro:2001:BCA

[51] Andrea Sgarro. Blind coding: authentication frauds from the point of view of rate-distortion theory. Journal of Discrete Mathematical Sciences and Cryptography, 4(2-3):133–150, 2001. CODEN ???? ISSN 0972-0529.

Kulli:2001:BGB

[52] V. R. Kulli and M. S. Biradar. The blict graph and blitact graph of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 4 (2-3):151–162, 2001. CODEN ???? ISSN 0972-0529.

Singh:2001:ITA

[53] Yogesh Singh and C. S. Rai. An information theoretic approach to blind source separation. Journal of Discrete Mathematical Sciences and Cryptography, 4(2-3):163–175, 2001. CODEN ???? ISSN 0972-0529.

Chang:2001:SSC

[54] Yanxun Chang, Giovanni Lo Faro, and Hao Shen. Support sizes of completely decomposable triple systems. Journal of Discrete Mathematical Sciences and Cryptography, 4(2-3):177–188, 2001. CODEN ???? ISSN 0972-0529.

Honkala:2001:PUE

[55] Iiro Honkala and Tero Laihonen. On the probability of undetected error for binary codes used for combined correction and detection. Journal of Discrete Mathematical Sciences and Cryptography, 4(2-3):189–197, 2001. CODEN ???? ISSN 0972-0529.

Maccari:2002:SER

[56] Alicia Maccari, Olga Rueda, and Vilma Viazzi. A survey on edge reconstruction of graphs. Journal of Discrete Mathematical Sciences and Cryptography, 5(1):1–11, 2002. CODEN ???? ISSN 0972-0529.

Antonopoulou:2002:UAP

[57] Ira Antonopoulou. A user authentication protocol based on the intractability of the 3-coloring problem. Journal of Discrete Mathematical Sciences and Cryptography, 5(1):17–21, 2002. CODEN ???? ISSN 0972-0529.

Chu:2002:RFN

[58] Ta-Chung Chu. Ranking fuzzy numbers by a ratio of relative areas. Journal of Discrete Mathematical Sciences and Cryptography, 5(1):23–34, 2002. CODEN ???? ISSN 0972-0529.

Eugeni:2002:KDM

[59] Franco Eugeni and Daniela Tondini. Klein’s discrete model: the case of odd order. Journal of Discrete Mathematical Sciences and Cryptography, 5(1):35–41, 2002. CODEN ???? ISSN 0972-0529.

Stangarone:2002:EFS

[60] Rosa Stangarone and Luisella Verdi. Essentially s-fold secure authentication systems of level l. Journal of Discrete Mathematical Sciences and Cryptography, 5(1):43–50, 2002. CODEN ???? ISSN 0972-0529.

Maturo:2002:RBS

[61] Antonio Maturo and Mauro Zannetti. Rédei blocking sets with two Rédei lines and quasigroups. Journal of Discrete Mathematical Sciences and Cryptography, 5(1):51–62, 2002. CODEN ???? ISSN 0972-0529.

Nath:2002:DRM

[62] Prem Nath. On determining a representation for measures of entropy possessing branching property without assuming symmetry. Journal of Discrete Mathematical Sciences and Cryptography, 5(1):63–75, 2002. CO-DEN ???? ISSN 0972-0529.

Yeh:2002:CMA

[63] Yi-Shiung Yeh, Chu-Hsing Lin, and Wei-Shen Lai. Construct message authentication code with SHA and AEA. Journal of Discrete Mathematical Sciences and Cryptography, 5(1):77–83, 2002. CODEN ???? ISSN 0972-0529.

Zahedi:2002:AFN

[64] M. M. Zahedi and H. Moradi Kouchi. (Anti-) fuzzy (normal) sub-F - polygroups and fuzzy α-cosets. Journal of Discrete Mathematical Sciences and Cryptography, 5(1):85–104, 2002. CODEN ???? ISSN 0972-0529.

Chi:2002:ESD

[65] Cynthia Chi, Daniela Ferrero, Charles Hallford, and Rebecca Williams. On the edge sums of de Bruijn graphs. Journal of Discrete Mathematical Sciences and Cryptography, 5(2):105–117, 2002. CODEN ???? ISSN 0972-0529.

Gentile:2002:HC

[66] Giuseppe Gentile and Renato Migliorato. Hypergroupoids and cryptosystems. Journal of Discrete Mathematical Sciences and Cryptography, 5(2): 119–138, 2002. CODEN ???? ISSN 0972-0529.

Ferri:2002:EQ

[67] Osvaldo Ferri, Stefania Ferri, and Antonella Giallonardo. On the elliptic quadrics of PG(2s + 1, q). Journal of Discrete Mathematical Sciences and Cryptography, 5(2):139–143, 2002. CODEN ???? ISSN 0972-0529.

Arumugam:2002:NGC

[68] S. Arumugam, Indra Rajasingh, and P. Roushini Leely Pushpam. A note on the graphoidal covering number of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 5(2):145–150, 2002. CODEN ???? ISSN 0972-0529.

Hwang:2002:IDS

[69] Min-Shiang Hwang, Chao-Chen Yang, and Shiang-Feng Tzeng. Improved digital signature scheme based on factoring and discrete logarithms. Journal of Discrete Mathematical Sciences and Cryptography, 5(2):151–155, 2002. CODEN ???? ISSN 0972-0529.

Girtis:2002:DTIa

[70] Konstantinos G. Girtis and George A. Tsihrintzis. Diffraction tomographic imaging algorithms. I. Data models and inversion. Journal of Discrete Mathematical Sciences and Cryptography, 5(2):157–178, 2002. CODEN ???? ISSN 0972-0529.

Chu:2002:GDA

[71] Y. P. Chu, W. C. Chiang, Y. H. Huang, and C. L. Lee. A generalized delay allocation model for rate-based packet scheduling algorithms. Journal of Discrete Mathematical Sciences and Cryptography, 5(2):179–197, 2002. CODEN ???? ISSN 0972-0529.

Chiu:2002:MPA

[72] Chin Ching Chiu and Yi-Shiung Yeh. A massively parallel algorithm based on DNA computations for constructing minimal spanning trees. Journal of Discrete Mathematical Sciences and Cryptography, 5(2):199–214, 2002. CODEN ???? ISSN 0972-0529.

Aggarwal:2002:CSN

[73] M. L. Aggarwal and Veena Budhraja. On construction of some new symmetric and asymmetric orthogonal arrays. Journal of Discrete Mathematical Sciences and Cryptography, 5(3):215–225, 2002. CODEN ???? ISSN 0972-0529.

Ferri:2002:STS

[74] Stefania Ferri. On some type (1, m, n)_{2} sets of PG(3, q). Journal of Discrete Mathematical Sciences and Cryptography, 5(3):227–230, 2002. CO-DEN ???? ISSN 0972-0529.

Mitton:2002:MNE

[75] Michel Mitton. On maximally nonlinear and extremal balanced Boolean functions. Journal of Discrete Mathematical Sciences and Cryptography, 5(3):231–253, 2002. CODEN ???? ISSN 0972-0529.

Foundas:2002:CIC

[76] E. Foundas and G. Chondrocoukis. Classifications and inferences of collective preferences. Journal of Discrete Mathematical Sciences and Cryptography, 5(3):255–280, 2002. CODEN ???? ISSN 0972-0529.

Girtis:2002:DTIb

[77] Konstantinos G. Girtis and George A. Tsihrintzis. Diffraction tomographic imaging algorithms. II. Random data inversion and super-resolution. Journal of Discrete Mathematical Sciences and Cryptography, 5(3):281–301, 2002. CODEN ???? ISSN 0972-0529.

Chen:2003:CDB

[78] T. W. Chen and J. Y. Lin. Control design based on minimization of quadratic index with an observer-based structure. Journal of Discrete Mathematical Sciences and Cryptography, 6(1):1–19, 2003. CODEN ???? ISSN 0972-0529.

Armeni:2003:SIH

[79] S. Armeni, D. Christodulakis, I. Kostopoulos, Y. C. Stamatiou, and M. Xenos. Secure information hiding based on computationally intractable problems. Journal of Discrete Mathematical Sciences and Cryptography, 6(1):21–33, 2003. CODEN ???? ISSN 0972-0529.

Aggarwal:2003:NBI

[80] M. L. Aggarwal and Seema Gupta. Nested balanced incomplete block designs derivable from simplex codes. Journal of Discrete Mathematical Sciences and Cryptography, 6(1):35–44, 2003. CODEN ???? ISSN 0972-0529.

Lin:2003:KUS

[81] Chu-Hsing Lin, Yi-Shiung Yeh, and Chen-Yu Lee. Keyed/unkeyed SHA-2. Journal of Discrete Mathematical Sciences and Cryptography, 6(1):45–58, 2003. CODEN ???? ISSN 0972-0529.

Bhandari:2003:CRS

[82] M. C. Bhandari and C. Durairajan. On the covering radius of simplex codes. Journal of Discrete Mathematical Sciences and Cryptography, 6 (1):59–69, 2003. CODEN ???? ISSN 0972-0529.

Ferri:2003:ICC

[83] Stefania Ferri. On the irreducible cubs containing the maximum number of points of a Fano plane. Journal of Discrete Mathematical Sciences and Cryptography, 6(1):71–75, 2003. CODEN ???? ISSN 0972-0529.

Zannetti:2003:CCQ

[84] Mauro Zannetti. A combinatorial characterization of quadric Veronesean. Journal of Discrete Mathematical Sciences and Cryptography, 6(1):77–82, 2003. CODEN ???? ISSN 0972-0529.

Marcu:2003:NBH

[85] Dănuţ Marcu. A note on the balanced hypergraphs. Journal of Discrete Mathematical Sciences and Cryptography, 6(1):83–84, 2003. CODEN ???? ISSN 0972-0529.

Kheddouci:2003:DRG

[86] Hamamache Kheddouci. A dual of 4-regular graph for G × C_{2n}. Journal of Discrete Mathematical Sciences and Cryptography, 6(1):85–96, 2003. CODEN ???? ISSN 0972-0529.

Borzooei:2003:SKP

[87] R. A. Borzooei, P. Corsini, and M. M. Zahedi. Some kinds of positive implicative hyper K-ideals. Journal of Discrete Mathematical Sciences and Cryptography, 6(1):97–108, 2003. CODEN ???? ISSN 0972-0529.

Ameri:2003:CHH

[88] R. Ameri. On categories of hypergroups and hypermodules. Journal of Discrete Mathematical Sciences and Cryptography, 6(2-3):121–132, 2003. CODEN ???? ISSN 0972-0529. Algebraic hyperstructures and applications.

Corsini:2003:CH

[89] Piergiulio Corsini. On Chinese hyperstructures. Journal of Discrete Mathematical Sciences and Cryptography, 6(2-3):133–137, 2003. CODEN ???? ISSN 0972-0529. Algebraic hyperstructures and applications.

Hort:2003:CHO

[90] Daniel Hort. A construction of hypergroups from ordered structures and their morphisms. Journal of Discrete Mathematical Sciences and Cryptography, 6(2-3):139–150, 2003. CODEN ???? ISSN 0972-0529. Algebraic hyperstructures and applications.

Gentile:2003:NQS

[91] Giuseppe Gentile. A note on quasi-Steiner hypergroupoids. Journal of Discrete Mathematical Sciences and Cryptography, 6(2-3):151–160, 2003. CODEN ???? ISSN 0972-0529. Algebraic hyperstructures and applications.

Iranmanesh:2003:THC

[92] A. Iranmanesh and A. H. Babareza. Transposition hypergroups and complement hypergroups. Journal of Discrete Mathematical Sciences and Cryptography, 6(2-3):161–168, 2003. CODEN ???? ISSN 0972-0529. Algebraic hyperstructures and applications.

Jantosciak:2003:SIF

[93] James Jantosciak and Christos G. Massouros. Strong identities and fortification in transposition hypergroups. Journal of Discrete Mathematical Sciences and Cryptography, 6(2-3):169–193, 2003. CODEN ???? ISSN 0972-0529. Algebraic hyperstructures and applications.

Leoreanu:2003:CHA

[94] Violeta Leoreanu. Commutative hypergroups associated with arbitrary hypergroups. Journal of Discrete Mathematical Sciences and Cryptography, 6(2-3):195–198, 2003. CODEN ???? ISSN 0972-0529. Algebraic hyperstructures and applications.

Madanshekaf:2003:SAH

[95] A. Madanshekaf. H_{v}-structures associated with P Q-hyperoperations. Journal of Discrete Mathematical Sciences and Cryptography, 6(2-3):199– 205, 2003. CODEN ???? ISSN 0972-0529. Algebraic hyperstructures and applications.

Massouros:2003:AHO

[96] Gerasimos G. Massouros. On the attached hypergroups of the order of an automation. Journal of Discrete Mathematical Sciences and Cryptography, 6(2-3):207–215, 2003. CODEN ???? ISSN 0972-0529. Algebraic hyperstructures and applications.

Ciampi:2003:PMH

[97] Rita Procesi Ciampi and Rosaria Rota. Polynomials over multiplicative hyperrings. Journal of Discrete Mathematical Sciences and Cryptography, 6(2-3):217–225, 2003. CODEN ???? ISSN 0972-0529. Algebraic hyperstructures and applications.

Barghi:2003:CH

[98] A. Rahnamai Barghi. A class of hyperrings. Journal of Discrete Mathematical Sciences and Cryptography, 6(2-3):227–233, 2003. CODEN ???? ISSN 0972-0529. Algebraic hyperstructures and applications.

Vougiouklis:2003:S

[99] Thomas Vougiouklis. The h/v-structures. Journal of Discrete Mathematical Sciences and Cryptography, 6(2-3):235–243, 2003. CODEN ???? ISSN 0972-0529. Algebraic hyperstructures and applications.

Migliorato:2003:CHH

[100] Renato Migliorato. Canonical v-hypergroups and ω-hypergroups. Journal of Discrete Mathematical Sciences and Cryptography, 6(2-3):245–256, 2003. CODEN ???? ISSN 0972-0529. Algebraic hyperstructures and applications.

Zannetti:2003:CCB

[101] Mauro Zannetti. A combinatorial characterization of Baer subspaces. Journal of Discrete Mathematical Sciences and Cryptography, 7(1):1–3, 2003. CODEN ???? ISSN 0972-0529.

Buccimazza:2003:HDB

[102] Biagio Buccimazza, Bal Kishan Dass, and Sapna Jain. High-density-burst error detection. Journal of Discrete Mathematical Sciences and Cryptography, 7(1):5–21, 2003. CODEN ???? ISSN 0972-0529.

Lu:2003:RKA

[103] Cheng-Fen Lu and Shiuh-Pyng Shieh. Re-keying algorithms for symmetric cryptosystems. Journal of Discrete Mathematical Sciences and Cryptography, 7(1):23–27, 2003. CODEN ???? ISSN 0972-0529.

DiGennaro:2003:WBS

[104] Fernando Di Gennaro. Which blocking sets are hypergroups? Journal of Discrete Mathematical Sciences and Cryptography, 7(1):29–36, 2003. CODEN ???? ISSN 0972-0529.

Chung:2003:DPG

[105] Hannah Chung, Daniela Ferrero, Alan Taylor, and Jeremy Warshauer. Diameter of path graphs. Journal of Discrete Mathematical Sciences and Cryptography, 7(1):49–54, 2003. CODEN ???? ISSN 0972-0529.

Chen:2003:BOT

[106] Wei-Kuei Chen and Chun-I Fan. Blind one-time signature and its applications. Journal of Discrete Mathematical Sciences and Cryptography, 7 (1):55–69, 2003. CODEN ???? ISSN 0972-0529.

Yeh:2003:CMS

[107] Yi-Shiung Yeh and Ta-Shan Tsui. A concurrent multi-string matching from multi-text algorithm based on the algorithm of Hamiltonian path problem and DNA computation. Journal of Discrete Mathematical Sciences and Cryptography, 7(1):71–96, 2003. CODEN ???? ISSN 0972-0529.

Selvakumar:2003:KCC

[108] R. Selvakumar and P. Balasubramani. Kernel code and its controllability. Journal of Discrete Mathematical Sciences and Cryptography, 7(1):97–101, 2003. CODEN ???? ISSN 0972-0529.

Yamuna:2003:EEF

[109] M. Yamuna and N. Sridharan. Every γ-excellent, γ-flexible graph is γ_{bi}- excellent. Journal of Discrete Mathematical Sciences and Cryptography, 7(1):103–110, 2003. CODEN ???? ISSN 0972-0529.

Hsu:2003:NCS

[110] Ching-Hung Hsu and Ming-Chih Lai. A new cipher scheme based on oneway hash function and IDEA. Journal of Discrete Mathematical Sciences and Cryptography, 7(1):111–120, 2003. CODEN ???? ISSN 0972-0529.

Tseng:2003:CKA

[111] Yuh-Min Tseng. A cryptographic key assignment scheme for access control in a hierarchy with exceptions. Journal of Discrete Mathematical Sciences and Cryptography, 7(1):121–132, 2003. CODEN ???? ISSN 0972-0529.

Girtis:2004:DTI

[112] Konstantinos G. Girtis and George A. Tsihrintzis. Diffraction tomographic imaging algorithms. III. Angularly limited data. Journal of Discrete Mathematical Sciences and Cryptography, 7(2):133–159, 2004. CODEN ???? ISSN 0972-0529.

Amato:2004:IPD

[113] Alberto Amato. Intersections and parallelism in P_{3}-designs. Journal of Discrete Mathematical Sciences and Cryptography, 7(2):161–166, 2004. CODEN ???? ISSN 0972-0529.

Dey:2004:EIL

[114] Partha P. Dey. Exploring invariant linear codes through generators and centralizers. Journal of Discrete Mathematical Sciences and Cryptography, 7(2):167–178, 2004. CODEN ???? ISSN 0972-0529.

Arumugam:2004:GUM

[115] S. Arumugam, Indra Rajasingh, and P. Roushini Leely Pushpam. Graphs with unique minimum acyclic graphoidal cover. I. Journal of Discrete Mathematical Sciences and Cryptography, 7(2):179–186, 2004. CODEN ???? ISSN 0972-0529.

Ferri:2004:SSI

[116] Stefania Ferri. On some singular irreducible cubics of PG(2, 2). Journal of Discrete Mathematical Sciences and Cryptography, 7(2):203–208, 2004. CODEN ???? ISSN 0972-0529.

Interlando:2004:ECI

[117] J. Carmelo Interlando and Michele Elia. On the effective computation of isomorphisms between σ_{F}(O_{F}/√) and GF(N_{F}(√)). Journal of Discrete

Mathematical Sciences and Cryptography, 7(2):209–225, 2004. CODEN ???? ISSN 0972-0529.

Mitton:2004:TUB

[118] Michel Mitton. Theoretical upper bounds on the covering radii of Boolean functions. Journal of Discrete Mathematical Sciences and Cryptography, 7(2):237–248, 2004. CODEN ???? ISSN 0972-0529.

Antonopoulou:2004:TP

[119] Hera Antonopoulou. On threshold properties. Journal of Discrete Mathematical Sciences and Cryptography, 7(2):249–254, 2004. CODEN ???? ISSN 0972-0529.

Rajola:2004:STS

[120] Sandro Rajola and Maria Scafati Tallini. The sets of type (0, n) and the Steiner triple systems in a finite affine plane. Journal of Discrete Mathematical Sciences and Cryptography, 7(3):261–266, 2004. CODEN ???? ISSN 0972-0529.

Ferri:2004:AST

[121] Stefania Ferri. Addendum to: “On some type (1, m, n)_{2} sets of PG(3, q)” [J. Discrete Math. Sci. Cryptography 5 (2002), no. 3, 227–230; Cno. 1986732]. Journal of Discrete Mathematical Sciences and Cryptography, 7(3):267–269, 2004. CODEN ???? ISSN 0972-0529.

Chu:2004:CDM

[122] Ta-Chung Chu, Kuo-Shi Huang, and Tzu-Ming Chang. COA defuzzification method for evaluating C_{pk} under fuzzy environments. Journal of Discrete Mathematical Sciences and Cryptography, 7(3):271–280, 2004. CODEN ???? ISSN 0972-0529.

Wu:2004:FSB

[123] Lin-Chuan Wu, Yi-Shiung Yeh, and Chun-I. Fan. Fail-stop blind signature scheme based on the integer factorization. Journal of Discrete Mathematical Sciences and Cryptography, 7(3):281–290, 2004. CODEN ???? ISSN 0972-0529.

Georgiou:2004:NVC

[124] S. Georgiou and C. Koukouvinos. New visual cryptographic schemes derived from orthogonal and mixed orthogonal arrays. Journal of Discrete Mathematical Sciences and Cryptography, 7(3):291–306, 2004. CODEN ???? ISSN 0972-0529.

Kotsireas:2004:IHM

[125] Ilias S. Kotsireas and Christos Koukouvinos. Inequivalent Hadamard matrices with buckets. Journal of Discrete Mathematical Sciences and Cryptography, 7(3):307–317, 2004. CODEN ???? ISSN 0972-0529.

Poulos:2004:CAC

[126] Marios Poulos, Maria Rangoussi, Vasilios Chrissikopoulos, Agelos Evangelou, and Fotis Georgiacodis. Comparative analysis of the computational geometry and neural network classification methods for person identification purposes via the EEG. I. Journal of Discrete Mathematical Sciences and Cryptography, 7(3):319–347, 2004. CODEN ???? ISSN 0972-0529.

Kofahi:2004:PSS

[127] Najib A. Kofahi, Turki Al-Somani, and Khalid Al-Zamil. Performance study of some symmetric block cipher algorithms under Linux operating system. Journal of Discrete Mathematical Sciences and Cryptography, 7 (3):359–370, 2004. CODEN ???? ISSN 0972-0529.

Wang:2005:VCA

[128] Shiuh-Jeng Wang. Vowel-consonant addressing mode on hashing for English letter-oriented keys. Journal of Discrete Mathematical Sciences and Cryptography, 8(1):1–25, 2005. CODEN ???? ISSN 0972-0529.

Al-Hassan:2005:ITM

[129] Qassem M. Al-Hassan. On inverses of tridiagonal matrices. Journal of Discrete Mathematical Sciences and Cryptography, 8(1):49–58, 2005. CO-DEN ???? ISSN 0972-0529.

Balasubramani:2005:CDP

[130] P. Balasubramani and R. Viswanathan. Characterizations of the 0distributive poset. Journal of Discrete Mathematical Sciences and Cryptography, 8(1):81–98, 2005. CODEN ???? ISSN 0972-0529.

Nath:2005:MSS

[131] Prem Nath. On the measurable solutions of some functional equations useful in information theory. Journal of Discrete Mathematical Sciences and Cryptography, 8(1):99–115, 2005. CODEN ???? ISSN 0972-0529.

Effantin:2005:EVC

[132] Brice Effantin and Hamamache Kheddouci. Exact values for the bchromatic number of a power complete k-ary tree. Journal of Discrete Mathematical Sciences and Cryptography, 8(1):117–129, 2005. CODEN ???? ISSN 0972-0529.

Geiselmann:2005:KSA

[133] Willi Geiselmann and Rainer Steinwandt. A key substitution attack on SFLASH^{v3}. Journal of Discrete Mathematical Sciences and Cryptography, 8(2):137–141, 2005. CODEN ???? ISSN 0972-0529.

Iembo:2005:CFP

[134] Rosanna Iembo, Massimo Squillante, and Aldo Ventre. Common fixed points for families of commutative transformations. Journal of Discrete Mathematical Sciences and Cryptography, 8(2):175–188, 2005. CODEN ???? ISSN 0972-0529.

Koukouvinos:2005:SCC

[135] C. Koukouvinos and P. Mantas. Some combinatorial constructions for optimal mixed-level supersaturated designs. Journal of Discrete Mathematical Sciences and Cryptography, 8(2):189–201, 2005. CODEN ???? ISSN 0972-0529.

Ratiner:2005:MBC

[136] Michael Ratiner. The method of S-box construction. Journal of Discrete Mathematical Sciences and Cryptography, 8(2):203–215, 2005. CODEN ???? ISSN 0972-0529.

Swaminathan:2005:SVM

[137] V. Swaminathan and P. Jeyanthi. On super vertex-magic labeling. Journal of Discrete Mathematical Sciences and Cryptography, 8(2):217–224, 2005. CODEN ???? ISSN 0972-0529.

Berardi:2005:BSB

[138] Luigia Berardi, Bal Kishan Dass, and Sapna Jain. Bound for solid burst error correction with Lee weight. Journal of Discrete Mathematical Sciences and Cryptography, 8(2):225–235, 2005. CODEN ???? ISSN 0972-0529.

Lucci:2005:CCT

[139] Maurizio Lucci. A combinatorial characterization of two skew lines. Journal of Discrete Mathematical Sciences and Cryptography, 8(2):237–239, 2005. CODEN ???? ISSN 0972-0529.

Kotsireas:2005:GAC

[140] Ilias S. Kotsireas and Christos Koukouvinos. Genetic algorithms for the construction of Hadamard matrices with two circulant cores. Journal of Discrete Mathematical Sciences and Cryptography, 8(2):241–250, 2005. CODEN ???? ISSN 0972-0529.

Yiming:2005:NPC

[141] Abulimiti Yiming and Masami Yasuda. A note on properties for a complementary graph and its tree graph. Journal of Discrete Mathematical Sciences and Cryptography, 8(2):251–259, 2005. CODEN ???? ISSN 0972-0529.

Ferri:2005:CC

[142] Stefania Ferri. The cubic curves of PG(2, 3). Journal of Discrete Mathematical Sciences and Cryptography, 8(2):261–266, 2005. CODEN ???? ISSN 0972-0529.

Lucci:2005:SC

[143] Maurizio Lucci. On (q + 1)^{2}-sets of class [0, 1, m, n]_{1} in PG(4, q). Journal of Discrete Mathematical Sciences and Cryptography, 8(2):267–270, 2005. CODEN ???? ISSN 0972-0529.

To:2005:SCS

[144] Vu Dong Tô, Reihaneh Safavi-Naini, and Yejing Wang. 2-secure codes from 2-SFP codes. Journal of Discrete Mathematical Sciences and Cryptography, 8(2):287–304, 2005. CODEN ???? ISSN 0972-0529.

Chang:2005:PSS

[145] Ming-Hsin Chang and Yi-Shiung Yeh. Proxy signature schemes with time limitation. Journal of Discrete Mathematical Sciences and Cryptography, 8(3):305–314, 2005. CODEN ???? ISSN 0972-0529.

Mitton:2005:AII

[146] Michel Mitton. An analysis and an improvement of iterative fast correlation attacks. Journal of Discrete Mathematical Sciences and Cryptography, 8(3):315–342, 2005. CODEN ???? ISSN 0972-0529.

Yeh:2005:REM

[147] Yi-Shiung Yeh, I-Te Chen, and Chan-Chi Wang. Randomizing encryption mode. Journal of Discrete Mathematical Sciences and Cryptography, 8(3): 355–364, 2005. CODEN ???? ISSN 0972-0529.

Salavati:2005:GP

[148] N. Shajareh-Pour Salavati. On the generalized permutations. Journal of Discrete Mathematical Sciences and Cryptography, 8(3):365–372, 2005. CODEN ???? ISSN 0972-0529.

Basavanagoud:2005:GEL

[149] B. Basavanagoud and Veena N. Mathad. Graph equations for line graphs, blitact graphs and blict graphs. Journal of Discrete Mathematical Sciences and Cryptography, 8(3):373–379, 2005. CODEN ???? ISSN 0972-0529.

Fan:2005:CL

[150] Chen-Ming Fan and H. J. Shyr. δ-codes and δ-languages. Journal of Discrete Mathematical Sciences and Cryptography, 8(3):381–394, 2005. CO-DEN ???? ISSN 0972-0529.

Tan:2005:CPT

[151] Ngo Dac Tan. On the classification problem for tetravalent metacirculant graphs. Journal of Discrete Mathematical Sciences and Cryptography, 8 (3):403–412, 2005. CODEN ???? ISSN 0972-0529.

Vijayarangan:2005:PPT

[152] N. Vijayarangan and R. Vijayasarathy. Primitive polynomials testing methodology. Journal of Discrete Mathematical Sciences and Cryptography, 8(3):427–435, 2005. CODEN ???? ISSN 0972-0529.

Lo:2005:OEC

[153] Chih-Yao Lo. The operation of the elliptic curve cryptology in IPv6. Journal of Discrete Mathematical Sciences and Cryptography, 8(3):437– 446, 2005. CODEN ???? ISSN 0972-0529.

Mascella:2005:CPB

[154] Raffaele Mascella and Luca G. Tallini. Combinatorial properties of balanced codes over the complex m-th roots of unity and efficient coding algorithms. Journal of Discrete Mathematical Sciences and Cryptography, 8(3):447–458, 2005. CODEN ???? ISSN 0972-0529.

Mascella:2005:EHO

[155] Raffaele Mascella and Luca G. Tallini. On efficient high-order spectral-null codes over the m-ary alphabet. Journal of Discrete Mathematical Sciences and Cryptography, 8(3):459–481, 2005. CODEN ???? ISSN 0972-0529.

Effantin:2005:CNP

[156] Brice Effantin. The b-chromatic number of power graphs of complete caterpillars. Journal of Discrete Mathematical Sciences and Cryptography, 8(3):483–502, 2005. CODEN ???? ISSN 0972-0529.

Suganya:2006:LRK

[157] A. Suganya and N. Vijayarangan. Landau–Ramanujan keyed hash functions for message authentication. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):1–8, 2006. CODEN ???? ISSN 0972-0529.

Chakrabarti:2006:NVC

[158] Sucheta Chakrabarti and R. K. Khanna. A new visual cryptography scheme for STS based access structures. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):9–23, 2006. CODEN ???? ISSN 0972-0529.

Elia:2006:RPS

[159] Michele Elia. Representation of primes as the sums of two squares in the golden section quadratic field. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):25–37, 2006. CODEN ???? ISSN 0972-0529.

Gaeta:2006:GAF

[160] Matteo Gaeta, Gerardo Iovane, and Enver Sangineto. A 3D geometric approach to face detection and facial expression recognition. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):39–53, 2006. CO-DEN ???? ISSN 0972-0529.

Wang:2006:SNG

[161] Jinhua Wang. The spectrum of nested group divisible designs of type t^{n}. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):55–65, 2006. CODEN ???? ISSN 0972-0529.

Sundaram:2006:PPC

[162] S. Soma Sundaram, A. Nagarajan, and M. Anantha Krishnan. Path point cover. Journal of Discrete Mathematical Sciences and Cryptography, 9(1): 67–71, 2006. CODEN ???? ISSN 0972-0529.

Mojdeh:2006:ERG

[163] Doost Ali Mojdeh. On the extending of k-regular graphs and their strong defining spectrum. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):73–86, 2006. CODEN ???? ISSN 0972-0529.

Chang:2006:STS

[164] Yung-Tsang Chang, Jung-Hua Hsu, and S. S. Yu. A sublinear time string prefix square detection algorithm. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):87–95, 2006. CODEN ???? ISSN 0972-0529.

Alderson:2006:SGR

[165] T. Alderson. Some generalizations of Rédei’s theorem. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):97–106, 2006. CODEN ???? ISSN 0972-0529.

Evangelaras:2006:NLC

[166] H. Evangelaras and C. Koukouvinos. New light on certain two level designs using Gröbner bases. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):107–124, 2006. CODEN ???? ISSN 0972-0529.

Evangelaras:2006:EAI

[167] H. Evangelaras, C. Koukouvinos, and E. Lappas. An efficient algorithm for the identification of isomorphic orthogonal arrays. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):125–132, 2006. CODEN ???? ISSN 0972-0529.

Chang:2006:AFE

[168] Chin-Chen Chang and Jung-San Lee. An anonymous and flexible t-out-of-n electronic voting scheme. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):133–151, 2006. CODEN ???? ISSN 0972-0529.

Yang:2006:WBG

[169] Cheng-Hsing Yang and Shiuh-Jeng Wang. Weighted bipartite graph for locating optimal LSB substitution for secret embedding. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):153–164, 2006. CO-DEN ???? ISSN 0972-0529.

Zhao:2006:CPP

[170] Yu-Kuang Zhao. A characterization of prefix n-power words. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):165–176, 2006. CODEN ???? ISSN 0972-0529.

Mojdeh:2006:CGG

[171] Doost Ali Mojdeh and Nader Jafari Rad. Connected geodomination in graphs. Journal of Discrete Mathematical Sciences and Cryptography, 9 (1):177–186, 2006. CODEN ???? ISSN 0972-0529.

Mascella:2006:TDA

[172] Raffaele Mascella and Luca G. Tallini. Theory and design of m-ary balanced codes which are invariant under symbol permutation. Journal of Discrete Mathematical Sciences and Cryptography, 9(1):187–200, 2006. CODEN ???? ISSN 0972-0529.

Zahedi:2006:CHA

[173] M. M. Zahedi and L. Torkzadeh. A classification of hyper K-algebras of order 3. Journal of Discrete Mathematical Sciences and Cryptography, 9 (2):201–213, 2006. CODEN ???? ISSN 0972-0529.

Karunagaram:2006:DPM

[174] Bala Shanthi Karunagaram and J. Paulraj Joseph. On domination parameters and maximum degree of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 9(2):215–223, 2006. CODEN ???? ISSN 0972-0529.

Ahrabian:2006:PGA

[175] H. Ahrabian and A. Nowzari-Dalini. Parallel generation of t-ary trees with ballot-sequences. Journal of Discrete Mathematical Sciences and Cryptography, 9(2):225–234, 2006. CODEN ???? ISSN 0972-0529.

Tan:2006:NCM

[176] Ngo Dac Tan and Chawalit Iamjaroen. A necessary condition for maximal nonhamiltonian Burkard–Hammer graphs. Journal of Discrete Mathematical Sciences and Cryptography, 9(2):235–252, 2006. CODEN ???? ISSN 0972-0529.

Sethumadhavan:2006:CAB

[177] M. Sethumadhavan, C. Yogha Lakxmie, and C. Vijayagovindan. A construction of p-ary balanced sequence with large k-error linear complexity. Journal of Discrete Mathematical Sciences and Cryptography, 9(2):253– 261, 2006. CODEN ???? ISSN 0972-0529.

Manuel:2006:LTN

[178] Paul Manuel, Bharati Rajan, Indra Rajasingh, and M. Chris Monica. Landmarks in torus networks. Journal of Discrete Mathematical Sciences and Cryptography, 9(2):263–271, 2006. CODEN ???? ISSN 0972-0529.

Ferri:2006:SAI

[179] Stefania Ferri. The singular absolutely irreducible cubics of PG(2, 3). Journal of Discrete Mathematical Sciences and Cryptography, 9(2):291– 305, 2006. CODEN ???? ISSN 0972-0529.

Hirasawa:2006:EEB

[180] Shigeichi Hirasawa and Masao Kasahara. Exponential error bounds and decoding complexity for block concatenated codes with tail biting trellis inner codes. Journal of Discrete Mathematical Sciences and Cryptography, 9(2):307–320, 2006. CODEN ???? ISSN 0972-0529.

Yeh:2006:DE

[181] Yi-Shiung Yeh, I-Te Chen, Ting-Yu Huang, and Chan-Chi Wang. Dynamic extended DES. Journal of Discrete Mathematical Sciences and Cryptography, 9(2):321–330, 2006. CODEN ???? ISSN 0972-0529.

Georgiou:2006:EDE

[182] S. Georgiou, C. Koukouvinos, and E. Lappas. Extremal doubly-even selfdual codes from Hadamard matrices. Journal of Discrete Mathematical Sciences and Cryptography, 9(2):331–339, 2006. CODEN ???? ISSN 0972-0529.

Evangelaras:2006:IDL

[183] H. Evangelaras and C. Koukouvinos. Interaction detection in Latin and hyper-Latin squares. Journal of Discrete Mathematical Sciences and Cryptography, 9(2):341–348, 2006. CODEN ???? ISSN 0972-0529.

Poinsot:2006:BFF

[184] Laurent Poinsot. Bent functions on a finite nonabelian group. Journal of Discrete Mathematical Sciences and Cryptography, 9(2):349–364, 2006. CODEN ???? ISSN 0972-0529.

Koukouvinos:2006:BSM

[185] C. Koukouvinos, P. Mantas, and S. Stylianou. A block-stepwise method for analyzing a specific type of supersaturated designs. Journal of Discrete Mathematical Sciences and Cryptography, 9(2):383–402, 2006. CODEN ???? ISSN 0972-0529.

Kinsley:2006:DBA

[186] A. Anto Kinsley and S. Somasundaram. Domination based algorithm to k-center problem. Journal of Discrete Mathematical Sciences and Cryptography, 9(3):403–416, 2006. CODEN ???? ISSN 0972-0529.

Mitton:2006:WFA

[187] Michel Mitton. On the Walsh–Fourier analysis of Boolean functions. Journal of Discrete Mathematical Sciences and Cryptography, 9(3):429–439, 2006. CODEN ???? ISSN 0972-0529.

Kung:2006:BML

[188] Ching-Jing Kung. 32-bit multipliers for linear congruential random number generators. Journal of Discrete Mathematical Sciences and Cryptography, 9(3):441–448, 2006. CODEN ???? ISSN 0972-0529.

Wang:2006:KMP

[189] Shiuh-Jeng Wang and Pei-Shun Hu. Key management for pay-TV broadcast systems in hierarchy. Journal of Discrete Mathematical Sciences and Cryptography, 9(3):449–476, 2006. CODEN ???? ISSN 0972-0529.

Sgarro:2006:CDM

[190] Andrea Sgarro and Luca Bortolussi. Codeword distinguishability in minimum diversity decoding. Journal of Discrete Mathematical Sciences and Cryptography, 9(3):487–502, 2006. CODEN ???? ISSN 0972-0529.

Evangelaras:2006:HE

[191] H. Evangelaras, C. Koukouvinos, and K. Mylona. On Hadamard embeddability. Journal of Discrete Mathematical Sciences and Cryptography, 9 (3):503–512, 2006. CODEN ???? ISSN 0972-0529.

Rao:2006:ABW

[192] Koduganti Venkata Rao and P. S. Avadhani. Authentication based on wavelet transformations. Journal of Discrete Mathematical Sciences and Cryptography, 9(3):513–521, 2006. CODEN ???? ISSN 0972-0529.

Chang:2006:EIB

[193] Chin-Chen Chang and Jen-Ho Yang. An efficient ID-based user identification scheme for mutual authentication in mobile environments. Journal of Discrete Mathematical Sciences and Cryptography, 9(3):537–548, 2006. CODEN ???? ISSN 0972-0529.

Chang:2006:CWA

[194] Ya-Fen Chang, Chin-Chen Chang, and Chia-Lin Kao. Comments on WLT attack on an enhanced timestamp-based password authentication scheme. Journal of Discrete Mathematical Sciences and Cryptography, 9(3):549– 555, 2006. CODEN ???? ISSN 0972-0529.

Butkote:2006:CSD

[195] R. Butkote and K. Denecke. Characterization of sets of discrete functions by algebraic identities. Journal of Discrete Mathematical Sciences and Cryptography, 9(3):557–572, 2006. CODEN ???? ISSN 0972-0529.

Chang:2006:SEY

[196] Chin-Chen Chang and Jung-San Lee. Security enhancement of Yang–Li hierarchical access control mechanism. Journal of Discrete Mathematical Sciences and Cryptography, 9(3):573–582, 2006. CODEN ???? ISSN 0972-0529.

Gan:2006:EMR

[197] C. S. Gan and V. C. Koo. Enumerations of the maximum rectilinear crossing numbers of complete and complete multi-partite graphs. Journal of Discrete Mathematical Sciences and Cryptography, 9(3):583–590, 2006. CODEN ???? ISSN 0972-0529.

Yeh:2007:ASM

[198] Yi-Shiung Yeh, Ting-Yu Huang, I-Te Chen, and Shih-Chin Chou. Analyze SHA-1 in message schedule. Journal of Discrete Mathematical Sciences and Cryptography, 10(1):1–7, February 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=01.

Quistorff:2007:FCT

[199] Jörn Quistorff. Fundamentals of coding type problems. Journal of Discrete Mathematical Sciences and Cryptography, 10(1):9–39, February 2007. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=10&issue_id=01.

Ameri:2007:HS

[200] R. Ameri and H. Hedayati. On k-hyperideals of semihyperrings. Journal of Discrete Mathematical Sciences and Cryptography, 10(1):41–54, 2007. CODEN ???? ISSN 0972-0529.

Ameri:2007:KHS

[201] R. Ameri and H. Hedayati. On k-hyperideals of semihyperrings. Journal of Discrete Mathematical Sciences and Cryptography, 10(1):41– 54, February 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=01.

Alsultanny:2007:SAI

[202] Yas Alsultanny. Systematic approach for image partitioning and shuffling. Journal of Discrete Mathematical Sciences and Cryptography, 10 (1):55–71, February 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=01.

Krishna:2007:GSD

[203] A. V. N. Krishna, S. N. N. Pandit, and A. Vinaya Babu. A generalized scheme for data encryption technique using a randomized matrix key. Journal of Discrete Mathematical Sciences and Cryptography, 10(1):73–81, February 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=01.

Barrera-Figueroa:2007:NAK

[204] V. Barrera-Figueroa, J. Sosa-Pedroza, and J. López-Bonilla. Numerical approach to King’s analytical study for circular loop antenna. Journal of Discrete Mathematical Sciences and Cryptography, 10(1): 83–92, February 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=01.

Chang:2007:RNI

[205] Ya-Fen F. Chang, Chin-Chen C. Chang, and Yi-Fang F. Cheng. Robust non-interactive t-out-of-n oblivious transfer based on the RSA cryptosystem. Journal of Discrete Mathematical Sciences and Cryptography, 10(1): 93–109, February 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=01.

Al-Ibrahim:2007:DVS

[206] Mohamed Hussain Al-Ibrahim. A designated verifier signature using secret sharing technique. Journal of Discrete Mathematical Sciences and Cryptography, 10(1):111–123, February 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=01.

Jahan:2007:RDT

[207] Israt Jahan and Mohammad Zahidur Rahman. A realistic divisible transferable electronic cash for general use. Journal of Discrete Mathematical Sciences and Cryptography, 10(1):125–150, February 2007. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=10&issue_id=01.

Borges-Quintana:2007:GBS

[208] M. Borges-Quintana, M. A. Borges-Trenard, and E. Martínez-Moro. On a Gröbner bases structure associated to linear codes. Journal of Discrete Mathematical Sciences and Cryptography, 10(2):151–191, April 2007. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=10&issue_id=02.

Wu:2007:APS

[209] Chia-Chi C. Wu and Chin-Chen C. Chang. Attacks on provably secure proxy-protected signature schemes based on factoring. Journal of Discrete Mathematical Sciences and Cryptography, 10(2):193–204, April 2007. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=10&issue_id=02.

Ratsaby:2007:DBF

[210] J. Ratsaby. Dimension and Boolean functions with long runs. Journal of Discrete Mathematical Sciences and Cryptography, 10(2):205– 225, April 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=02.

Ratsaby:2007:VDB

[211] Joel Ratsaby. On the VC-dimension and Boolean functions with long runs. Journal of Discrete Mathematical Sciences and Cryptography, 10 (2):205–225, 2007. CODEN ???? ISSN 0972-0529.

Jen:2007:SMC

[212] Hsiao-Lih L. Jen, Tai-Chang C. Hsia, and M. Babul Hasan. A study of methods for construction of balanced incomplete block design. Journal of Discrete Mathematical Sciences and Cryptography, 10(2):227– 243, April 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=02.

Santhakumaran:2007:PRC

[213] A. P. Santhakumaran. Periphery with respect to cliques in graphs. Journal of Discrete Mathematical Sciences and Cryptography, 10(2): 245–254, April 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=02.

Georgiou:2007:SDC

[214] S. Georgiou, C. Koukouvinos, and E. Lappas. Self-dual codes over some prime fields constructed from skew-Hadamard matrices. Journal of Discrete Mathematical Sciences and Cryptography, 10(2):255–266, April 2007.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=10&issue_id=02.

Mukhopadhyay:2007:FBA

[215] Debdeep Mukhopadhyay and Dipanwita RoyChowdhury. Fault based attack of the Rijndael cryptosystem. Journal of Discrete Mathematical Sciences and Cryptography, 10(2):267–290, April 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=02.

Kumari:2007:IMC

[216] Meena Kumari. Indices to measure the cryptographic strength of Sbox. Journal of Discrete Mathematical Sciences and Cryptography, 10 (2):291–312, April 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=02.

Antonopoulou:2007:AEB

[217] Sotiria Antonopoulou, Yannis C. Stamatiou, and Malvina Vamvakari. An asymptotic expansion for the q-binomial series using singularity analysis for generating functions. Journal of Discrete Mathematical Sciences and Cryptography, 10(3):313–328, 2007. CODEN ???? ISSN 0972-0529.

Antonopoulou:2007:AEQ

[218] S. Antonopoulou, Y. C. Stamatiou, and M. Vamvakari. An asymptotic expansion for the q-binomial series using singularity analysis for generating functions. Journal of Discrete Mathematical Sciences and Cryptography, 10(3):313–328, June 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=03.

Thiyagarajan:2007:VCC

[219] M. Thiyagarajan and N. Sairam. Verification of the correctness of compiler optimization using co-induction. Journal of Discrete Mathematical Sciences and Cryptography, 10(3):329–349, June 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=03.

Lee:2007:SEK

[220] Jung-San S. Lee and Chin-Chen C. Chang. Security enhancement of key authentication schemes based on discrete logarithms. Journal of Discrete Mathematical Sciences and Cryptography, 10(3):351–358, June 2007. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=10&issue_id=03.

Wang:2007:IHO

[221] S. J. Wang. Information hiding at oblivious watermarking scheme upon alternative threshold in spatial domains. Journal of Discrete Mathematical Sciences and Cryptography, 10(3):359–384, June 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=03.

Longani:2007:AFN

[222] V. Longani. An algorithm for finding the number of shortest routes on square lattices. Journal of Discrete Mathematical Sciences and Cryptography, 10(3):385–393, June 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=03.

Longani:2007:API

[223] Vites Longani. An algorithm for partitioning an integer n as a sum of k positive integers. Journal of Discrete Mathematical Sciences and Cryptography, 10(3):395–399, June 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=03.

Jormakka:2007:FFP

[224] Jorma Jormakka. On finding Fermat’s pairs. Journal of Discrete Mathematical Sciences and Cryptography, 10(3):401–413, June 2007. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=10&issue_id=03.

Santhakumaran:2007:EGN

[225] A. P. Santhakumaran and J. John. Edge geodetic number of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 10(3): 415–432, June 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=03.

Maralabhavi:2007:LES

[226] Y. B. Maralabhavi and Venkanagouda M. Goudar. Lict edge semientire graph of a planar graph. Journal of Discrete Mathematical Sciences and Cryptography, 10(3):433–438, June 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=03.

Laccetti:2007:BFA

[227] G. Laccetti and G. Schmid. Brute force attacks on hash functions. Journal of Discrete Mathematical Sciences and Cryptography, 10(3): 439–460, June 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=03.

Manuel:2007:VLB

[228] Paul Manuel, Kalim Qureshi, Albert William, and Albert Muthumalai. VLSI layout of Benes networks. Journal of Discrete Mathematical Sciences and Cryptography, 10(4):461–472, August 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=04.

Baril:2007:HPI

[229] Jean-Lue L. Baril. Hamiltonian paths for involutions in the square of a Cayley graph. Journal of Discrete Mathematical Sciences and Cryptography, 10(4):473–484, August 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=04.

Kari:2007:ICR

[230] Lila Kari and Kalpana Mahalingam. k-involution codes and related sets. Journal of Discrete Mathematical Sciences and Cryptography, 10(4):485– 503, 2007. CODEN ???? ISSN 0972-0529.

Kari:2007:KIC

[231] L. Kari and K. Mahalingam. k-involution codes and related sets. Journal of Discrete Mathematical Sciences and Cryptography, 10(4):485– 503, August 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=04.

Chaluvaraju:2007:CTD

[232] B. Chaluvaraju and N. D. Soner. Complementary total domination in graphs. Journal of Discrete Mathematical Sciences and Cryptography, 10 (4):505–516, August 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=04.

Shanthini:2007:UFG

[233] A. P. Shanthini, A. Antonysamy, and T. Tamil Chelvam. Userfriendly genetic algorithm. Journal of Discrete Mathematical Sciences and Cryptography, 10(4):517–536, August 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=04.

Ibrahim:2007:CSS

[234] Aminu A. Ibrahim. A counting scheme and some algebraic properties of a class of special permutation patterns. Journal of Discrete Mathematical Sciences and Cryptography, 10(4):537–546, August 2007. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=10&issue_id=04.

Lee:2007:EMR

[235] C. C. Lee, M. S. Hwang, and C. W. Lin. An efficient multi-round anonymous auction protocol. Journal of Discrete Mathematical Sciences and Cryptography, 10(4):547–557, August 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=04.

Deacon:2007:PKT

[236] Stephanie Deacon, Eduardo Dueñez, and José Iovino. A public-key threshold crypto system based on residue rings. Journal of Discrete Mathematical Sciences and Cryptography, 10(4):559–571, August 2007. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=10&issue_id=04.

Lin:2007:AAA

[237] Robert H.-J. Lin, Chih-Pin Yen, Peter S. Deng, K.-C. Fan, Jennifer Shu-Jen Lin, and Peter Chu. Analysis of the ϵ-algorithm to accelerate the convergence. Journal of Discrete Mathematical Sciences and Cryptography, 10(4):573–580, August 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=04.

Dayantis:2007:SPL

[238] Jean Dayantis and Jean-François F. Palierne. A search for primes from lesser primes. Journal of Discrete Mathematical Sciences and Cryptography, 10(4):581–602, August 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=04.

Chang:2007:MAK

[239] Chin-Chen C. Chang, Chia-Yin Y. Lee, and Yu-Wei W. Su. A mutual authenticated key agreement scheme over insecure networks. Journal of Discrete Mathematical Sciences and Cryptography, 10(5):603– 612, October 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=05.

Kulli:2007:ITD

[240] V. R. Kulli and R. R. Iyer. Inverse total domination in graphs. Journal of Discrete Mathematical Sciences and Cryptography, 10(5):613– 620, October 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=05.

Ikpotokin:2007:MDM

[241] F. O. Ikpotokin and S. C. Chiemeke. Mathematical derivation of the multi-peg Tower of Hanoi algorithm. Journal of Discrete Mathematical Sciences and Cryptography, 10(5):621–636, October 2007. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=10&issue_id=05.

Darafsheh:2007:CSO

[242] M. R. Darafsheh and Y. Farjami. Calculating the set of orders of elements in the finite linear groups. Journal of Discrete Mathematical Sciences and Cryptography, 10(5):637–653, October 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=05.

Chen:2007:SOP

[243] Hon-Chan C. Chen. A simple optimal parallel algorithm for constructing a spanning tree of a trapezoid graph. Journal of Discrete Mathematical Sciences and Cryptography, 10(5):655–660, October 2007. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=10&issue_id=05.

Prasad:2007:PDN

[244] B. Jaya Prasad, T. Tamizh Chelvam, and S. Robinson Chellathurai. Private domination number of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 10(5):661–666, October 2007. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=10&issue_id=05.

Chelvam:2007:CPD

[245] T. Tamizh Chelvam and B. Jaya Prasad. Common point domination number. Journal of Discrete Mathematical Sciences and Cryptography, 10(5): 667–676, October 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=05.

Santhakumaran:2007:ESN

[246] A. P. Santhakumaran and J. John. The edge Steiner number of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 10 (5):677–696, October 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=05.

Reddy:2007:AKE

[247] P. Vasudeva Reddy and M. Padmavathamma. An authenticated key exchange protocol in elliptic curve cryptography. Journal of Discrete Mathematical Sciences and Cryptography, 10(5):697–705, October 2007. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=10&issue_id=05.

Ogwu:2007:PTI

[248] F. J. Ogwu, M. Talib, and G. A. Aderounmu. The performance of TCP/IP over wireless ATM network. Journal of Discrete Mathematical Sciences and Cryptography, 10(5):707–720, October 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=05.

Sweetly:2007:RNG

[249] R. Sweetly and J. Paulraj Joseph. Radio number of graphs of diameter three. Journal of Discrete Mathematical Sciences and Cryptography, 10(5):721–730, October 2007. CODEN ???? ISSN 0972-0529.

URL http://www.connectjournals.com/achivestoc.php?bookmark=
CJ-003072&volume=10&issue_id=05.

Sathyanarayana:2007:GPS

[250] S. V. Sathyanarayana, M. Aswatha Kumar, and K. N. Hari Bhat. Generation of pseudorandom sequence over elliptic curve group and their properties. Journal of Discrete Mathematical Sciences and Cryptography, 10(6):731–747, December 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=06.

Hsu:2007:CAD

[251] K. S. Hsu, R. C. Soong, K. C. Chen, and T. S. Lan. A computerized approach to the design of automobile suspension system. Journal of Discrete Mathematical Sciences and Cryptography, 10(6):749–761, December 2007.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=10&issue_id=06.

Berardi:2007:SBS

[252] Luigia Berardi and Manuela Romani. On the size of a blocking set in a symmetric design. Journal of Discrete Mathematical Sciences and Cryptography, 10(6):763–774, December 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=06.

Ferri:2007:RST

[253] Osvaldo Ferri and Stefania Ferri. A remark on the Segrè theorem. Journal of Discrete Mathematical Sciences and Cryptography, 10(6):775– 778, December 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=06.

Saxena:2007:IBG

[254] Jyoti Saxena, C. S. Rai, and P. K. Bansal. ICA based GSIC receiver for multirate DS-CDMA systems. Journal of Discrete Mathematical Sciences and Cryptography, 10(6):779–792, December 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=06.

Poinsot:2007:PNB

[255] Laurent Poinsot. Perfect nonlinear S-boxes on the real-line. Journal of Discrete Mathematical Sciences and Cryptography, 10(6):793– 813, December 2007. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=10&issue_id=06.

Seenivasan:2007:TML

[256] M. Seenivasan, A. Lourdusamy, and M. Ravi Ramasubramanian. Triangular mean labeling of graphs. Journal of Discrete Mathematical Sciences and Cryptography, 10(6):815–822, December 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=06.

Ferri:2007:SGC

[257] Stefania Ferri. On some graphic characterizations of cubic curves in PG(2, 2) and in PG(2, 3). Journal of Discrete Mathematical Sciences and Cryptography, 10(6):823–843, December 2007. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=10&issue_id=06.

Anonymous:2007:IV

[258] Anonymous. Index of volume 10. Journal of Discrete Mathematical Sciences and Cryptography, 10(6):845–850, December 2007. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=10&issue_id=06.

Babu:2008:PCR

[259] N. S. Babu. On perfect codes in rank distance. Journal of Discrete Mathematical Sciences and Cryptography, 11(1):1–11, February 2008. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=11&issue_id=01.

Marcu:2008:GRR

[260] D. Marcu. A generalization of a result of A. Renyi. Journal of Discrete Mathematical Sciences and Cryptography, 11(1):13–17, February 2008.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=11&issue_id=01.

Marcu:2008:SKO

[261] D. Marcu. A study on the k-ordered Hamiltonian graphs. Journal of Discrete Mathematical Sciences and Cryptography, 11(1):19– 31, February 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=01.

Huang:2008:WAU

[262] Y. F. Huang and K. L. Wen. Weighting analysis by using grey method. Journal of Discrete Mathematical Sciences and Cryptography, 11(1):33– 58, February 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=01.

Om:2008:ATC

[263] H. Om and R. Patwa. Affine transformation in cryptography. Journal of Discrete Mathematical Sciences and Cryptography, 11(1):59– 65, February 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=01.

Ghiani:2008:LCP

[264] G. Ghiani, E. Manni, and C. Triki. The lane covering problem with time windows. Journal of Discrete Mathematical Sciences and Cryptography, 11(1):67–81, February 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=01.

Iovane:2008:LLB

[265] G. Iovane and E. Sangineto. LICoS: Landmark-based Identikit Composition and Suspect Retrieval. Journal of Discrete Mathematical Sciences and Cryptography, 11(1):83–98, February 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=01.

Hou:2008:DAD

[266] C. I. Hou, C. Y. Lo, and J. H. Leu. A digital application of data encryption. Journal of Discrete Mathematical Sciences and Cryptography, 11(1):99–111, February 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=01.

Foundas:2008:IAC

[267] E. Foundas, Ch. Lytras, and C. Patsakis. Improved algorithms for the calculation of Fibonacci numbers. Journal of Discrete Mathematical Sciences and Cryptography, 11(1):113–120, February 2008. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=11&issue_id=01.

Lin:2008:FAM

[268] R. H. J. Lin, G. Jung, J. S. J. Lin, and P. Chu. Further analysis on matrix operation of AHP. Journal of Discrete Mathematical Sciences and Cryptography, 11(2):121–130, April 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=02.

Hong-Song:2008:NNB

[269] C. Hong-Song, Z. Jianyu, and H. W. J. Lee. A novel NP-based security scheme for AODV routing protocol. Journal of Discrete Mathematical Sciences and Cryptography, 11(2):131–145, April 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=02.

Juarna:2008:CIB

[270] A. Juarna and V. Vajnovszki. Combinatorial isomorphism between Fibonacci classes. Journal of Discrete Mathematical Sciences and Cryptography, 11(2):147–158, April 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=02.

Kanso:2008:MNS

[271] A. A. Kanso. Mixer — a new stream cipher. Journal of Discrete Mathematical Sciences and Cryptography, 11(2):159–179, April 2008. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=11&issue_id=02.

Foundas:2008:LPS

[272] E. Foundas, C. Patsakis, and G. Chondrocoukis. Lucas permutations and some notes on Fibonacci permutations. Journal of Discrete Mathematical Sciences and Cryptography, 11(2):181–190, April 2008. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=11&issue_id=02.

Mitton:2008:BFV

[273] M. Mitton. Boolean functions of n variables and permutations on F_{n}. Journal of Discrete Mathematical Sciences and Cryptography, 11(2): 191–200, April 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=02.

Mitton:2008:MPC

[274] M. Mitton. Minimal polynomial of Cayley graph adjacency matrix for Boolean functions. Journal of Discrete Mathematical Sciences and Cryptography, 11(2):201–207, April 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=02.

Artikis:2008:RSI

[275] P. T. Artikis and C. T. Artikis. Random sums of integral part models in computer systems operations. Journal of Discrete Mathematical Sciences and Cryptography, 11(2):209–217, April 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=02.

Sundaram:2008:GLP

[276] M. Sundaram, R. Ponraj, and S. Somasundaram. On graph labeling parameters. Journal of Discrete Mathematical Sciences and Cryptography, 11(2):219–229, April 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=02.

Viswanath:2008:IVH

[277] M. K. Viswanath and A. R. Deepti. An improvised version of Hill’s cipher. Journal of Discrete Mathematical Sciences and Cryptography, 11 (2):231–237, April 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=02.

Ponraj:2008:MLG

[278] R. Ponraj and S. Somasundaram. Mean labeling of graphs obtained by identifying two graphs. Journal of Discrete Mathematical Sciences and Cryptography, 11(2):239–252, April 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=02.

Huang:2008:UCP

[279] T. L. Huang, C. W. Liu, C. C. Chao, K. T. Lee, T. Y. Hwang, and C. M. Chung. Upper core point detection using improved ant colony optimization algorithm. Journal of Discrete Mathematical Sciences and Cryptography, 11(3):253–265, June 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=03.

Islam:2008:PER

[280] M. I. Islam, M. Zulhasnine, and M. R. Amin. Performance evaluation of rake receiver of DS-CDMA under AWGN environments. Journal of Discrete Mathematical Sciences and Cryptography, 11(3):267–279, June 2008.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=11&issue_id=03.

Kim:2008:PFS

[281] Y. S. Kim and J. H. Chang. A proxy fail-stop signature scheme with proxy revocation. Journal of Discrete Mathematical Sciences and Cryptography, 11(3):281–295, June 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=03.

Sethumadhavan:2008:AKE

[282] M. Sethumadhavan, M. Sindhu, C. Srinivasan, and C. Kavitha. An algorithm for k-error joint linear complexity of binary multisequences. Journal of Discrete Mathematical Sciences and Cryptography, 11(3): 297–304, June 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=03.

Ibrahim:2008:CDC

[283] H. Ibrahim, R. Ramli, and M. H. Hassan. Combinatorial design for a conference: constructing a balanced three-parallel session schedule. Journal of Discrete Mathematical Sciences and Cryptography, 11(3): 305–317, June 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=03.

Kumar:2008:SSI

[284] Y. Kumar and N. R. Pillai. On shift sequences for interleaved construction of signal sets with low correlation. Journal of Discrete Mathematical Sciences and Cryptography, 11(3):319–331, June 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=03.

Das:2008:USL

[285] A. K. Das. An unconditionally secure location-aware key management scheme for static sensor networks. Journal of Discrete Mathematical Sciences and Cryptography, 11(3):333–355, June 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=03.

Chakrabarti:2008:NOE

[286] S. Chakrabarti and S. K. Pal. A new online extended visual cryptography scheme for STS access structures. Journal of Discrete Mathematical Sciences and Cryptography, 11(3):357–374, June 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=03.

Yeh:2008:OS

[287] Y. S. Yeh, M. C. Kao, and W. S. Lai. The orthogonal signature. Journal of Discrete Mathematical Sciences and Cryptography, 11(4): 375–384, August 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=04.

Wang:2008:LIL

[288] C. C. Wang, M. C. Kao, and Y. S. Yeh. Low information leakage random padding scheme for block encryption. Journal of Discrete Mathematical Sciences and Cryptography, 11(4):385–391, August 2008. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=11&issue_id=04.

Verma:2008:DCN

[289] A. K. Verma, M. Dave, and R. C. Joshi. DNA cryptography: a novel paradigm for secure routing in Mobile Ad hoc Networks (MANETs). Journal of Discrete Mathematical Sciences and Cryptography, 11(4): 393–404, August 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=04.

Glinos:2008:EBR

[290] N. Glinos and Y. C. Stamatiou. On the equivalence between random graph models. Journal of Discrete Mathematical Sciences and Cryptography, 11(4):405–419, August 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=04.

Jesintha:2008:NCG

[291] J. J. Jesintha and G. Sethuraman. A new class of graceful rooted trees. Journal of Discrete Mathematical Sciences and Cryptography, 11 (4):421–435, August 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=04.

Atani:2008:PKC

[292] R. E. Atani, S. E. Atani, and S. Mirzakuchaki. Public key cryptography using semigroup actions and semirings. Journal of Discrete Mathematical Sciences and Cryptography, 11(4):437–445, August 2008. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=11&issue_id=04.

Procesi:2008:MHR

[293] R. Procesi and R. Rota. Multiplicative HV-rings and complementary hyperstructures. Journal of Discrete Mathematical Sciences and Cryptography, 11(4):447–456, August 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=04.

Ferri:2008:GCT

[294] S. Ferri. Graphic characterizations of twisted cubic in PG(3, q). Journal of Discrete Mathematical Sciences and Cryptography, 11(4):457– 460, August 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=04.

Marcu:2008:ACN

[295] D. Marcu. An approximate chromatic number of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 11(4):461– 464, August 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=04.

Ikorong:2008:RGC

[296] G. Ikorong. A reformulation of the Goldbach conjecture. Journal of Discrete Mathematical Sciences and Cryptography, 11(4):465– 469, August 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=04.

Mathew:2008:JBH

[297] S. Mathew and K. P. Jacob. JERIM-320: a 320-bit hash function with dual applications — a statistical analysis. Journal of Discrete Mathematical Sciences and Cryptography, 11(4):471–482, August 2008. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=11&issue_id=04.

Clark:2008:ADR

[298] L. Clark. Ascents and descents in random trees. Journal of Discrete Mathematical Sciences and Cryptography, 11(4):483–492, August 2008.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=11&issue_id=04.

Basu:2008:DFO

[299] M. Basu, Md. M. Rahaman, and S. Bagchi. 3 − (2m, 2m|1, 2m|2|1) design for first order Reed–Muller code R(1, m). Journal of Discrete Mathematical Sciences and Cryptography, 11(4):493–500, August 2008. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=11&issue_id=04.

Han:2008:GPG

[300] N. Han and Z. Liang. On the graceful permutation graphs conjecture. Journal of Discrete Mathematical Sciences and Cryptography, 11 (5):501–526, October 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=05.

Azanchiler:2008:ELS

[301] H. Azanchiler. Extension of line-splitting operation from graphs to binary matroids. Journal of Discrete Mathematical Sciences and Cryptography, 11(5):527–536, October 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=05.

Khadir:2008:AFS

[302] O. Khadir. Algorithm for factoring some RSA and Rabin moduli. Journal of Discrete Mathematical Sciences and Cryptography, 11(5):537– 543, October 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=05.

Lee:2008:ED

[303] C. Y. Lee and W. S. Lai. Extended DSA. Journal of Discrete Mathematical Sciences and Cryptography, 11(5):545–550, October 2008. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=11&issue_id=05.

Ferri:2008:NSQ

[304] S. Ferri. On the non singular quartics of PG(2, 3). Journal of Discrete Mathematical Sciences and Cryptography, 11(5):551–558, October 2008.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=11&issue_id=05.

Germina:2008:FRS

[305] K. A. Germina, K. Abhishek, and K. L. Princy. Further results on set-valued graphs. Journal of Discrete Mathematical Sciences and Cryptography, 11(5):559–566, October 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=05.

Huang:2008:SAP

[306] P. H. Huang. A study of applying parallel THBL algorithm to a fighter plane system. Journal of Discrete Mathematical Sciences and Cryptography, 11(5):567–577, October 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=05.

Dhenakaran:2008:MSK

[307] S. S. Dhenakaran, E. R. Naganathan, and C. Ganesamoorthy. Multiple symmetric keys using Banach fixed point theorem. Journal of Discrete Mathematical Sciences and Cryptography, 11(5):579–587, October 2008.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=11&issue_id=05.

Chandra:2008:GMB

[308] P. Chandra, R. K. Singh, and Y. Singh. Generation of mutants for Boolean expression. Journal of Discrete Mathematical Sciences and Cryptography, 11(5):589–607, October 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=05.

Chang:2008:MDA

[309] H. T. Chang. Mathematical development and analysis to conflict resolution. Journal of Discrete Mathematical Sciences and Cryptography, 11(5): 609–618, October 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=05.

Anonymous:2008:EVN

[310] Anonymous. Errata: Volume 11 number 5. Journal of Discrete Mathematical Sciences and Cryptography, 11(5):619, October 2008. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=11&issue_id=05.

Dey:2008:NCP

[311] P. P. Dey and M. Asifuzzaman. A note on the code of projective plane of order four. Journal of Discrete Mathematical Sciences and Cryptography, 11(6):621–637, December 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=06.

Guo:2008:MIM

[312] F. Guo, Y. Mu, and Z. Chen. Multi-identity management for identitybased cryptography. Journal of Discrete Mathematical Sciences and Cryptography, 11(6):639–672, December 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=06.

Sharma:2008:IEA

[313] M. Sharma, M. K. Kowar, and M. Sharma. An improved evolutionary algorithm for secured image using adaptive genetic algorithm. Journal of Discrete Mathematical Sciences and Cryptography, 11(6):673– 683, December 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=06.

Cinque:2008:CSL

[314] L. Cinque, G. Iovane, and E. Sangineto. Comparing SIFT and LDA based face recognition approaches. Journal of Discrete Mathematical Sciences and Cryptography, 11(6):685–704, December 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=06.

Mou:2008:DCA

[315] S. Mou, Md. I. Islam, and M. R. Amin. Dynamic channel allocation in mobile cellular networks. Journal of Discrete Mathematical Sciences and Cryptography, 11(6):705–714, December 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=06.

Nemron:2008:OAT

[316] Ikorong Anouk Gilbert Nemron. An original abstract over the twin primes, the Goldbach conjecture, the friendly numbers, the perfect numbers, the Mersenne composite numbers, and the Sophie Germain primes. Journal of Discrete Mathematical Sciences and Cryptography, 11(6):715–726, December 2008. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=11&issue_id=06.

Seenivasan:2008:VEL

[317] M. Seenivasan and A. Lourdusamy. Vertex equitable labeling of graphs. Journal of Discrete Mathematical Sciences and Cryptography, 11(6):727– 735, December 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=11&issue_id=06.

Navaei:2008:HMH

[318] L. Navaei and R. Dayanian. Hidden message via hypothesis testing and information — theoretic model. Journal of Discrete Mathematical Sciences and Cryptography, 11(6):737–746, December 2008. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=11&issue_id=06.

Anonymous:2008:IV

[319] Anonymous. Index of volume 11 (2008). Journal of Discrete Mathematical Sciences and Cryptography, 11(6):747–752, December 2008. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=11&issue_id=06.

Iovane:2009:MTU

[320] G. Iovane, G. Ingenito, and M. Leone. Motion tracking using fuzzy logic and consistent labeling for multiple objects in multiple cameras vision. Journal of Discrete Mathematical Sciences and Cryptography, 12(1):1–42, February 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=01.

Kazemi:2009:GGI

[321] A. P. Kazemi. Geodomination in graphs II. Journal of Discrete Mathematical Sciences and Cryptography, 12(1):43–50, February 2009. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=12&issue_id=01.

Islam:2009:PCT

[322] Md. I. Islam and J. K. Das. Performance comparison of two and three beam SDMA traffic in mobile cellular network. Journal of Discrete Mathematical Sciences and Cryptography, 12(1):51–61, February 2009. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=12&issue_id=01.

Sitthiwiratham:2009:PJG

[323] T. Sitthiwiratham and C. Promsakon. Planarity of joined graphs. Journal of Discrete Mathematical Sciences and Cryptography, 12(1): 63–69, February 2009. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=12&issue_id=01.

Longani:2009:APT

[324] V. Longani. Applications of Polya’s theorem to distribution problems and partitions of integers. Journal of Discrete Mathematical Sciences and Cryptography, 12(1):71–77, February 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=01.

Al-Hassan:2009:IPM

[325] Q. Al-Hassan. On inverses of pentadiagonal matrices. Journal of Discrete Mathematical Sciences and Cryptography, 12(1):79–91, February 2009.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=12&issue_id=01.

Ghameshlou:2009:STD

[326] A. N. Ghameshlou and S. M. Sheikholeslami. Signed (total) domination numbers and Laplacian spectrum of graphs. Journal of Discrete Mathematical Sciences and Cryptography, 12(1):93–99, February 2009. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=12&issue_id=01.

Settepanella:2009:BSC

[327] S. Settepanella. Blocking sets in the complement of hyperplane arrangements in projective space. Journal of Discrete Mathematical Sciences and Cryptography, 12(1):101–107, February 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=01.

Lourdusamy:2009:GPN

[328] A. Lourdusamy and C. Muthulakshmi. Generalized t-pebbling number of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 12(1):109–120, February 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=01.

Nitaj:2009:AEC

[329] A. Nitaj. Application of ECM to a class of RSA keys. Journal of Discrete Mathematical Sciences and Cryptography, 12(2):121–137, April 2009. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=12&issue_id=02.

Saha:2009:PSK

[330] M. Saha and D. Roychowdhury. Provably secure key establishment protocol using one-way functions. Journal of Discrete Mathematical Sciences and Cryptography, 12(2):139–158, April 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=02.

Lizzio:2009:MEE

[331] A. Lizzio and M. A. Munda. Maximum embedding of a E2(v − w, 4, 1) into a P(v, 4, e). Journal of Discrete Mathematical Sciences and Cryptography, 12(2):159–166, April 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=02.

Sattanathan:2009:DPG

[332] M. Sattanathan and R. Kala. Degree prime graph. Journal of Discrete Mathematical Sciences and Cryptography, 12(2):167–173, April 2009. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=12&issue_id=02.

Cengellenmis:2009:GIC

[333] Y. Cengellenmis. The gray image of the cyclic codes over IF 2[u]/u_{2} − 1. Journal of Discrete Mathematical Sciences and Cryptography, 12(2): 175–178, April 2009. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=12&issue_id=02.

Chelvam:2009:NSD

[334] T. T. Chelvam and S. R. Chellathurai. A note on split domination number of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 12(2):179–186, April 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=02.

Ferri:2009:CMN

[335] S. Ferri. On class [m, n]1k-sets in PG(2, 3). Journal of Discrete Mathematical Sciences and Cryptography, 12(2):187–191, April 2009. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=12&issue_id=02.

Pawar:2009:SER

[336] Y. S. Pawar and P. V. Patil. Semiprime elements and relative annihilators in multiplicative lattices. Journal of Discrete Mathematical Sciences and Cryptography, 12(2):193–204, April 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=02.

Artikis:2009:IPM

[337] C. T. Artikis and P. T. Artikis. Integral part models in information systems and educational processes. Journal of Discrete Mathematical Sciences and Cryptography, 12(2):205–215, April 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=02.

Wu:2009:MAD

[338] C. C. Wu, Y. L. Wu, and P. C. Yang. Multiple attribute decision making assessment on the notebook computer industry performance. Journal of Discrete Mathematical Sciences and Cryptography, 12(2): 217–237, April 2009. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=12&issue_id=02.

Cengellenmis:2009:UCC

[339] Y. Cengellenmis. On (1 − u_{2})-cyclic and cyclic codes over IFp + uIFp + u_{2}IFp. Journal of Discrete Mathematical Sciences and Cryptography, 12 (2):239–243, April 2009. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=12&issue_id=02.

Elumalai:2009:GUC

[340] A. Elumalai and G. Sethuraman. Gracefulness of union of cycle with parallel chords and complete bipartite graphs or paths. Journal of Discrete Mathematical Sciences and Cryptography, 12(2):245–252, April 2009. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=12&issue_id=02.

Rozario:2009:PEC

[341] L. J. Rozario and Md. I. Islam. Performance estimation of call admission control schemes based voice/data integrated wireless cellular network with customer retrials. Journal of Discrete Mathematical Sciences and Cryptography, 12(3):253–294, June 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=03.

Basavanagoud:2009:PG

[342] B. Basavanagoud and K. G. Mirajkar. On permanent graphs. Journal of Discrete Mathematical Sciences and Cryptography, 12(3):295– 311, June 2009. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=12&issue_id=03.

Ismail:2009:NSS

[343] E. S. Ismail, N. M. F. Tahat, and R. R. Ahmad. A new signature scheme based on factoring and discrete logarithms. Journal of Discrete Mathematical Sciences and Cryptography, 12(3):313–318, June 2009. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=12&issue_id=03.

Rad:2009:NVC

[344] N. J. Rad. A note on yL/t-vertex critical graphs. Journal of Discrete Mathematical Sciences and Cryptography, 12(3):319–322, June 2009. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=12&issue_id=03.

Mitton:2009:BBF

[345] M. Mitton. On Boolean bent functions. Journal of Discrete Mathematical Sciences and Cryptography, 12(3):323–333, June 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=03.

Artikis:2009:DRS

[346] C. T. Artikis and P. T. Artikis. Discrete random sums in processes of systemics arising in cybernetics and informatics. Journal of Discrete Mathematical Sciences and Cryptography, 12(3):335–345, June 2009. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=12&issue_id=03.

Liang:2009:DCS

[347] Z. Liang and S. Yang. On DRC covering — a survey. Journal of Discrete Mathematical Sciences and Cryptography, 12(3):347–351, June 2009. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=12&issue_id=03.

Padmavathy:2009:PAI

[348] R. Padmavathy and C. Bhagvati. Performance analysis of index calculus method. Journal of Discrete Mathematical Sciences and Cryptography, 12(3):353–371, June 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=03.

Kala:2009:RBN

[349] R. Kala and T. R. N. Vasantha. Restrained bondage number of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 12 (3):373–380, June 2009. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=12&issue_id=03.

Liang:2009:COP

[350] Z. Liang and J. Wang. Constructions of optimal packing and covering of the complete multigraph with applications. Journal of Discrete Mathematical Sciences and Cryptography, 12(4):381–410, August 2009. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=12&issue_id=04.

Alikhani:2009:DSC

[351] S. Alikhani and Y. H. Peng. Dominating sets of centipedes. Journal of Discrete Mathematical Sciences and Cryptography, 12(4):411– 428, August 2009. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=12&issue_id=04.

Hutsell:2009:EAR

[352] S. R. Hutsell and G. W. Greenwood. Efficient algebraic representation of quantum circuits. Journal of Discrete Mathematical Sciences and Cryptography, 12(4):429–449, August 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=04.

Om:2009:NNB

[353] H. Om and T. K. Sarkar. Neural network based intrusion detection system for detecting changes in hardware profile. Journal of Discrete Mathematical Sciences and Cryptography, 12(4):451–466, August 2009. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=12&issue_id=04.

Vilfred:2009:ISS

[354] V. Vilfred and C. Jayasekaran. Interchange similar self vertex switchings in graphs. Journal of Discrete Mathematical Sciences and Cryptography, 12(4):467–480, August 2009. CODEN ???? ISSN 0972-0529.

URL http://www.connectjournals.com/achivestoc.php?bookmark=
CJ-003072&volume=12&issue_id=04.

Acharya:2009:ENC

[355] B. D. Acharya, K. A. Germina, K. L. Princy, and S. B. Rao. Embedding and NP-completeness theorems for set-graceful, topologically setgraceful and set-sequential graphs. Journal of Discrete Mathematical Sciences and Cryptography, 12(4):481–487, August 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=04.

Kazemi:2009:DNS

[356] A. P. Kazemi. On the domination number of some graphs. Journal of Discrete Mathematical Sciences and Cryptography, 12(4):489– 493, August 2009. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=12&issue_id=04.

Lekshmi:2009:PMD

[357] R. S. Lekshmi and R. Anitha. Perfect matching dominating cycle — a new dominating cycle. Journal of Discrete Mathematical Sciences and Cryptography, 12(4):495–508, August 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=04.

Freyre:2009:SAR

[358] P. Freyre, N. Diaz, and E. R. Morgado. Some algorithms related to matrices with entries in a finite field. Journal of Discrete Mathematical Sciences and Cryptography, 12(5):509–519, October 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=05.

Hedayati:2009:CBI

[359] H. Hedayati and Z. Jafari. Connections between interval valued fuzzy graphs and fuzzy groups with (S, T)-norms. Journal of Discrete Mathematical Sciences and Cryptography, 12(5):521–531, October 2009. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=12&issue_id=05.

Freyre:2009:FAM

[360] P. Freyre, N. Diaz, and E. R. Morgado. Fast algorithm for the multiplication of a row vector by a randomly selected matrix A. Journal of Discrete Mathematical Sciences and Cryptography, 12(5):533–549, October 2009.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=12&issue_id=05.

Dayantis:2009:CRS

[361] J. Dayantis and J.-F. Palierne. Complementary results on the search of primes from lesser primes. Journal of Discrete Mathematical Sciences and Cryptography, 12(5):551–559, October 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=05.

Nemron:2009:SDP

[362] Ikorong Anouk Gilbert Nemron. A simplification of the difficult part of the Berge last conjecture. Journal of Discrete Mathematical Sciences and Cryptography, 12(5):561–569, October 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=05.

Nemron:2009:SDA

[363] Ikorong Anouk Gilbert Nemron. A simple dissertation around the Berge problem and the Hadwiger conjecture. Journal of Discrete Mathematical Sciences and Cryptography, 12(5):571–593, October 2009. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=12&issue_id=05.

Lee:2009:LPG

[364] C. C. Lee, K. Y. Huang, and S. Y. Huang. On-line password guessing attack on Lu–Cao key agreement protocol for secure authentication. Journal of Discrete Mathematical Sciences and Cryptography, 12 (5):595–598, October 2009. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=12&issue_id=05.

Daili:2009:SPA

[365] N. Daili. Some probabilistic approaches to some asymptotic prime numbers formulae. Journal of Discrete Mathematical Sciences and Cryptography, 12(5):599–614, October 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=05.

Koukouvinos:2009:ESU

[366] C. Koukouvinos, E. Lappas, and D. E. Simos. Encryption schemes using orthogonal arrays. Journal of Discrete Mathematical Sciences and Cryptography, 12(5):615–628, October 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=05.

Helmy:2009:UCO

[367] G. W. Helmy, M. A. Ahmed, and M. N. El-Derini. Using co-operating agents for partial results protection in mobile agent paradigm. Journal of Discrete Mathematical Sciences and Cryptography, 12(6):629– 647, December 2009. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=12&issue_id=06.

Chiu:2009:DTT

[368] M. C. Chiu, L. J. Yeh, and C. H. Yeh. The development of tactile testing machine on tact switch. Journal of Discrete Mathematical Sciences and Cryptography, 12(6):649–669, December 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=06.

Rani:2009:NSS

[369] R. U. Rani. A novel single source shortest path algorithm. Journal of Discrete Mathematical Sciences and Cryptography, 12(6):671– 680, December 2009. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=12&issue_id=06.

Talebi:2009:KIG

[370] Y. Talebi, M. S. Esmaeilifar, and S. Azizpour. A kind of intersection graph of vector space. Journal of Discrete Mathematical Sciences and Cryptography, 12(6):681–689, December 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=06.

Islam:2009:CAS

[371] Md. I. Islam, J. K. Das, and M. R. Amin. Call admission scheme of mixed traffic for mobile cellular networks. Journal of Discrete Mathematical Sciences and Cryptography, 12(6):691–705, December 2009. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=12&issue_id=06.

Zannetti:2009:CCP

[372] M. Zannetti. A combinatorial characterization of parabolic quadrics. Journal of Discrete Mathematical Sciences and Cryptography, 12(6):707– 715, December 2009. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=12&issue_id=06.

Huang:2009:PSC

[373] J. C. Huang and W. T. Pan. Performing stock closing price prediction through the use of principle component regression in association with general regression neural network. Journal of Discrete Mathematical Sciences and Cryptography, 12(6):717–728, December 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=06.

Sweetly:2009:RNW

[374] R. Sweetly and J. Paulraj Joseph. The radio number of (W_{n}: 2) graphs. Journal of Discrete Mathematical Sciences and Cryptography, 12(6):729– 736, December 2009. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=12&issue_id=06.

Stangarone:2009:CIO

[375] R. Stangarone and L. Verdi. Cryptosystems involving one factorizations of graphs. Journal of Discrete Mathematical Sciences and Cryptography, 12(6):737–743, December 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=06.

Poulakis:2009:PKE

[376] D. Poulakis. A public key encryption scheme based on factoring and discrete logarithm. Journal of Discrete Mathematical Sciences and Cryptography, 12(6):745–752, December 2009. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=12&issue_id=06.

Anonymous:2009:IV

[377] Anonymous. Index of volume 12 (2009). Journal of Discrete Mathematical Sciences and Cryptography, 12(6):753–758, December 2009. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=12&issue_id=06.

Anonymous:2009:CVN

[378] Anonymous. Contents — volume 12 number 6 (2009). Journal of Discrete Mathematical Sciences and Cryptography, 12(6):A, December 2009.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=12&issue_id=06.

Bahar:2010:FSI

[379] H. B. Bahar, M. Sokouti, and B. Sokouti. A first study of improving transposition cryptosystem. Journal of Discrete Mathematical Sciences and Cryptography, 13(1):1–9, February 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=01.

Ferri:2010:CCC

[380] S. Ferri. Combinatorial characterizations of (2q + 1)-caps in PG(r, q). Journal of Discrete Mathematical Sciences and Cryptography, 13(1):11– 15, February 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=01.

Lin:2010:DAM

[381] K. Y. Lin, J. C. Lin, J. M. Chen, and T. C. Wu. Defense automatic malicious tools based on navigation behavior. Journal of Discrete Mathematical Sciences and Cryptography, 13(1):17–27, February 2010. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=13&issue_id=01.

Padmavathy:2010:EKR

[382] R. Padmavathy and C. Bhagvati. Ephemeral key recovery using index calculus method. Journal of Discrete Mathematical Sciences and Cryptography, 13(1):29–43, February 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=01.

Mulaudzi:2010:EAP

[383] M. S. Mulaudzi and O. Akinyemi. q-enumeration of alternating permutations of odd length. Journal of Discrete Mathematical Sciences and Cryptography, 13(1):45–67, February 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=01.

Mitton:2010:CGD

[384] M. Mitton. Convolution and generalized difference sets. Journal of Discrete Mathematical Sciences and Cryptography, 13(1):69– 83, February 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=01.

Kobayashi:2010:ADE

[385] N. Kobayashi, T. Matsushima, and S. Hirasawa. An accurate density evolution analysis for a finite-state Markov channel. Journal of Discrete Mathematical Sciences and Cryptography, 13(1):85–97, February 2010.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=13&issue_id=01.

Shobhalatha:2010:LPD

[386] G. Shobhalatha and K. Suresh. Lattice properties in dioids. Journal of Discrete Mathematical Sciences and Cryptography, 13(1):99– 104, February 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=01.

Hirasawa:2010:NPG

[387] S. Hirasawa and M. Kasahara. A note on performance of generalized tail biting trellis codes. Journal of Discrete Mathematical Sciences and Cryptography, 13(2):105–122, April 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=02.

Hedayati:2010:IBF

[388] H. Hedayati. t-implication-based fuzzy interior hyperideals of semihypergroups. Journal of Discrete Mathematical Sciences and Cryptography, 13 (2):123–140, April 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=02.

Mohapatra:2010:WEP

[389] A. K. Mohapatra and N. Prakash. Wired equivalent privacy reinvestigated. Journal of Discrete Mathematical Sciences and Cryptography, 13 (2):141–151, April 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=02.

Cinque:2010:FRU

[390] L. Cinque, G. Iovane, M. Manzo, and E. Sangineto. Face recognition using SIFT features and a region-based ranking. Journal of Discrete Mathematical Sciences and Cryptography, 13(2):153–170, April 2010. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=13&issue_id=02.

Berardi:2010:CKE

[391] L. Berardi and R. Rota. A construction of a key exchange protocol. Journal of Discrete Mathematical Sciences and Cryptography, 13 (2):171–174, April 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=02.

Sharma:2010:EDM

[392] B. D. Sharma and N. Sookoo. Eigenvalues of the difference matrices of the Lee partition. Journal of Discrete Mathematical Sciences and Cryptography, 13(2):175–183, April 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=02.

Kouadri:2010:RBF

[393] A. Kouadri, L. Chiter, and M. Zelmat. A radial basis function neural network optimized through modified DIRECT algorithm based-model for a three interconnected water tank. Journal of Discrete Mathematical Sciences and Cryptography, 13(2):185–200, April 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=02.

Anonymous:2010:TCV

[394] Anonymous. Table of contents — volume 13 number 2. Journal of Discrete Mathematical Sciences and Cryptography, 13(2):A, April 2010. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=13&issue_id=02.

Feng:2010:FQH

[395] Y. Feng. p-fuzzy quasi-hypergroups obtained from fuzzy binary relations. Journal of Discrete Mathematical Sciences and Cryptography, 13 (3):201–208, June 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=03.

Porkodi:2010:SAE

[396] C. Porkodi, R. Arumuganathan, and K. Vidya. Single authority electronic voting based on elliptic curves. Journal of Discrete Mathematical Sciences and Cryptography, 13(3):209–217, June 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=03.

Hwang:2010:ENI

[397] S. J. Hwang and C. H. Chao. An efficient non-interactive deniable authentication protocol with anonymous sender protection. Journal of Discrete Mathematical Sciences and Cryptography, 13(3):219–231, June 2010. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=13&issue_id=03.

Pal:2010:DSC

[398] S. K. Pal, S. Kapoor, A. Arora, R. Chaudhary, and J. Khurana. Design of strong cryptographic schemes based on Latin Squares. Journal of Discrete Mathematical Sciences and Cryptography, 13(3):233–256, June 2010. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=13&issue_id=03.

Saxena:2010:SNC

[399] B. Saxena and S. K. Pal. Some new concepts in fuzzy arithmetic. Journal of Discrete Mathematical Sciences and Cryptography, 13(3): 257–270, June 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=03.

Li:2010:ELE

[400] R. Li. Eigenvalues, Laplacian eigenvalues, and Hamiltonian connectivity of graphs. Journal of Discrete Mathematical Sciences and Cryptography, 13(3):271–275, June 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=03.

Vasuki:2010:CNC

[401] R. Vasuki and A. Nagarajan. On the construction of new classes of super mean graphs. Journal of Discrete Mathematical Sciences and Cryptography, 13(3):277–290, June 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=03.

Fan:2010:NSM

[402] C. M. Fan and C. C. Huang. A note of solutions to the morphic involution language equations. Journal of Discrete Mathematical Sciences and Cryptography, 13(3):291–298, June 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=03.

Lee:2010:ALK

[403] M. J. Lee, M. K. Kuo, and Y. R. Chen. An antimagic labeling in K2n, 2n|nC4. Journal of Discrete Mathematical Sciences and Cryptography, 13(3):299–304, June 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=03.

Anonymous:2010:CVN

[404] Anonymous. Contents — volume 13 number 03 June 2010. Journal of Discrete Mathematical Sciences and Cryptography, 13(3):A, June 2010.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=13&issue_id=03.

Swapna:2010:BCI

[405] S. Swapna, A. D. Dileep, C. C. Sekhar, and S. Kant. Block cipher identification using support vector classification and regression. Journal of Discrete Mathematical Sciences and Cryptography, 13(4):305– 318, August 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=04.

Nagireddy:2010:IEM

[406] S. Nagireddy, H. A. Murthy, and S. Kant. Identification of encryption method for block ciphers using histogram method. Journal of Discrete Mathematical Sciences and Cryptography, 13(4):319–328, August 2010.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=13&issue_id=04.

Kant:2010:ISR

[407] S. Kant, V. Sharma, N. Verma, and B. K. Dass. Identification scheme for romanized Indian languages from their plain and ciphered bit stream. Journal of Discrete Mathematical Sciences and Cryptography, 13(4): 329–345, August 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=04.

Kant:2010:CPC

[408] S. Kant, R. K. Asthana, and B. L. Jangid. Comparative performance of the classifiers for cryptosystem identification. Journal of Discrete Mathematical Sciences and Cryptography, 13(4):347–368, August 2010. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=13&issue_id=04.

Saito:2010:NAC

[409] T. Saito, T. Matsushima, and S. Hirasawa. A note on automatic construction algorithms for orthogonal designs of experiments using error correcting codes. Journal of Discrete Mathematical Sciences and Cryptography, 13(4):369–381, August 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=04.

Suko:2010:APU

[410] T. Suko, T. Matsushima, and S. Hirasawa. Asymptotic property of universal lossless coding for independent piecewise identically distributed sources. Journal of Discrete Mathematical Sciences and Cryptography, 13(4):383–391, August 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=04.

Tang:2010:OMR

[411] Hui-Chin Tang, K. H. Hsieh, and Hwapeng Chang. A 4217th-order multiple recursive random number generator with modulus 2^{31} − 69. Journal of Discrete Mathematical Sciences and Cryptography, 13(4): 393–398, August 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=04.

Anonymous:2010:BR

[412] Anonymous. Book reviews. Journal of Discrete Mathematical Sciences and Cryptography, 13(4):399–405, August 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=04.

Anonymous:2010:CVA

[413] Anonymous. Contents — volume 13 (04 August 2010). Journal of Discrete Mathematical Sciences and Cryptography, 13(4):A, August 2010. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=13&issue_id=04.

Pal:2010:CAC

[414] S. Pal and P. Bhowmick. Cubic approximation of curve-shaped objects in Z_{2}: a generalized approach based on discrete curvature. Journal of Discrete Mathematical Sciences and Cryptography, 13(5):407– 427, October 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=05.

Po-Hsian:2010:WBD

[415] H. Po-Hsian. A weight-based data hiding method for binary image using selective binary numeral system. Journal of Discrete Mathematical Sciences and Cryptography, 13(5):429–444, October 2010. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=13&issue_id=05.

Vijayarangan:2010:HPE

[416] N. Vijayarangan. Hard problems in elliptic curve scalar multiplication. Journal of Discrete Mathematical Sciences and Cryptography, 13 (5):445–452, October 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=05.

Acharya:2010:VGG

[417] B. D. Acharya and K. A. Germina. Vertex-graceful graphs. Journal of Discrete Mathematical Sciences and Cryptography, 13(5):453– 463, October 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=05.

Daili:2010:ABD

[418] N. Daili. About binomial densities: some applications in number theory. Journal of Discrete Mathematical Sciences and Cryptography, 13(5): 465–472, October 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=05.

Ferri:2010:GCK

[419] O. Ferri and S. Ferri. Graphic characterizations of the k-caps and k-arcs in PG(r, q). Journal of Discrete Mathematical Sciences and Cryptography, 13(5):473–477, October 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=05.

Tang:2010:BLC

[420] Hui-Chin Tang, Kuang-Hang Hsieh, and Hwapeng Chang. A 32-bit linear congruential random number generator with prime modulus. Journal of Discrete Mathematical Sciences and Cryptography, 13(5):479– 486, October 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=05.

Reddy:2010:IBD

[421] P. V. Reddy, B. U. Rao, and T. Gowri. ID-based directed proxy signature scheme from bilinear pairings. Journal of Discrete Mathematical Sciences and Cryptography, 13(5):487–500, October 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=05.

Anonymous:2010:CVO

[422] Anonymous. Contents — volume 13 number 05 October 2010. Journal of Discrete Mathematical Sciences and Cryptography, 13(5):A, October 2010.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=13&issue_id=05.

Iovane:2010:BWW

[423] G. Iovane, P. Giordano, G. Ingenito, M. Leone, and A. Nocerino. Biometric watermarking via wavelet approach, reverse engineering in infosecurity and dematerialization of public administration. Journal of Discrete Mathematical Sciences and Cryptography, 13(6):501–542, December 2010.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=13&issue_id=06.

Kures:2010:CPS

[424] M. Kures. The composition of polynomials by the substitution principle. Journal of Discrete Mathematical Sciences and Cryptography, 13(6): 543–552, December 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=06.

Acharya:2010:SIG

[425] B. D. Acharya. Signed intersection graphs. Journal of Discrete Mathematical Sciences and Cryptography, 13(6):553–569, December 2010. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=13&issue_id=06.

Poinsot:2010:DPN

[426] L. Poinsot. Doubly perfect nonlinear Boolean permutations. Journal of Discrete Mathematical Sciences and Cryptography, 13(6):571– 582, December 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=06.

Kazemi:2010:EKK

[427] A. P. Kazemi. Every K1, 7 and K1, 3-free, 3-vertex critical graph of even order has a perfect matching. Journal of Discrete Mathematical Sciences and Cryptography, 13(6):583–591, December 2010. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=13&issue_id=06.

Tang:2010:BHA

[428] Hui-Chin Tang, K. H. Hsieh, and T. L. Chao. A backward heuristic algorithm for two-term multiple recursive random number generators. Journal of Discrete Mathematical Sciences and Cryptography, 13(6):593– 600, December 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=06.

Basavanagoud:2010:CNP

[429] B. Basavanagoud and P. V. Patil. A criterion for (non-)planarity of the transformation graph G(xyz) when xyz = − + +. Journal of Discrete Mathematical Sciences and Cryptography, 13(6):601–610, December 2010.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=13&issue_id=06.

Anonymous:2010:IV

[430] Anonymous. Index of volume 13 (2010). Journal of Discrete Mathematical Sciences and Cryptography, 13(6):611–616, December 2010. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=13&issue_id=06.

Anonymous:2010:CVD

[431] Anonymous. Contents — volume 13 number 06 December 2010. Journal of Discrete Mathematical Sciences and Cryptography, 13(6): A, December 2010. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=13&issue_id=06.

Chelvam:2011:NTG

[432] T. T. Chelvam and T. Asir. A note on total graph of Z(n). Journal of Discrete Mathematical Sciences and Cryptography, 14(1):1–7, February 2011.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=14&issue_id=01.

Al-Towaiq:2011:EAS

[433] M. H. Al-Towaiq and H. K. Al-Bzoor. An efficient algorithm for the solution of second order PDE’s using Taylor expansion polynomials. Journal of Discrete Mathematical Sciences and Cryptography, 14(1):9– 31, February 2011. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=14&issue_id=01.

Jaiyeola:2011:TCI

[434] T. G. Jaíyéo̧lá. On three cryptographic identities in left universal Osborn loops. Journal of Discrete Mathematical Sciences and Cryptography, 14(1):33–50, February 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=01.

Liang:2011:CPC

[435] Z. Liang. Coverings and packings of complete directed graphs by odd circuits. Journal of Discrete Mathematical Sciences and Cryptography, 14 (1):51–75, February 2011. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=14&issue_id=01.

Begum:2011:FDB

[436] N. Begum, M. Alam, and Md. I. Islam. Fingerprint detection based on DWT a new approach. Journal of Discrete Mathematical Sciences and Cryptography, 14(1):77–87, February 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=01.

Peter:2011:IAL

[437] S. J. Peter and S. P. Victor. An integrated approach for local outlier detection using dynamic minimum spanning tree. Journal of Discrete Mathematical Sciences and Cryptography, 14(1):89–106, February 2011.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=14&issue_id=01.

Anonymous:2011:CVF

[438] Anonymous. Contents — Vol. 14 # 01 February 2011. Journal of Discrete Mathematical Sciences and Cryptography, 14(1):A, February 2011.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=14&issue_id=01.

Elsheh:2011:CPK

[439] E. Elsheh and A. Youssef. Cryptanalysis of a public key cryptosystem based on boolean permutations. Journal of Discrete Mathematical Sciences and Cryptography, 14(2):107–115, April 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=02.

Chin:2011:SPI

[440] C. L. Chin and C. C. Huang. Some properties of involution codes. Journal of Discrete Mathematical Sciences and Cryptography, 14(2): 117–127, April 2011. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=14&issue_id=02.

Syahrul:2011:CTA

[441] E. Syahrul, J. Dubois, and V. Vajnovszki. Combinatorial transforms: Applications in lossless image compression. Journal of Discrete Mathematical Sciences and Cryptography, 14(2):129–147, April 2011. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=14&issue_id=02.

Peter:2011:MST

[442] S. J. Peter. Minimum spanning tree based clustering for outlier detection. Journal of Discrete Mathematical Sciences and Cryptography, 14 (2):149–166, April 2011. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=14&issue_id=02.

Peter:2011:SSC

[443] S. J. Peter. Structural similarity clustering for detection of outliers and hubs using minimum spanning tree. Journal of Discrete Mathematical Sciences and Cryptography, 14(2):167–183, April 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=02.

Kao:2011:EAT

[444] T. C. Kao and H. C. Tang. An exhaustive analysis of two-term multiple recursive random number generators with double precision floating point restricted multipliers. Journal of Discrete Mathematical Sciences and Cryptography, 14(2):185–191, April 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=02.

Vasuki:2011:FRS

[445] R. Vasuki and A. Nagarajan. Further results on super mean graphs. Journal of Discrete Mathematical Sciences and Cryptography, 14(2): 193–206, April 2011. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=14&issue_id=02.

Anonymous:2011:CVAa

[446] Anonymous. Contents — Vol. 14 # 02 April 2011. Journal of Discrete Mathematical Sciences and Cryptography, 14(2):A, April 2011. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=14&issue_id=02.

Iovane:2011:EAK

[447] G. Iovane, L. Puccio, G. Lamponi, and A. Amorosia. Electronic access key based on innovative Information Fusion technique involving prime numbers and biometric data. Journal of Discrete Mathematical Sciences and Cryptography, 14(3):207–225, June 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=03.

Peter:2011:SSM

[448] S. J. Peter. Structural similarity micro clustering algorithm for local outliers and hubs using dynamic minimum spanning tree. Journal of Discrete Mathematical Sciences and Cryptography, 14(3):227–247, June 2011. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=14&issue_id=03.

John:2011:FEC

[449] J. John, A. Vijayan, and S. Sujitha. The forcing edge covering number of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 14(3):249–259, June 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=03.

Bhaumik:2011:NIN

[450] J. Bhaumik and D. Roy Chowdhury. Nimix: An involutary nonlinear vectorial boolean function. Journal of Discrete Mathematical Sciences and Cryptography, 14(3):261–277, June 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=03.

Jebitha:2011:DTG

[451] M. K. A. Jebitha and J. P. Joseph. Domination in transformation graph G − +−. Journal of Discrete Mathematical Sciences and Cryptography, 14(3):279–303, June 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=03.

Anonymous:2011:CVJ

[452] Anonymous. Contents — Vol. 14 # 03 June 2011. Journal of Discrete Mathematical Sciences and Cryptography, 14(3):A, June 2011. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=14&issue_id=03.

Abdelhady:2011:NBE

[453] G. A. M. Abdelhady, F. S. Helail, and A. A. Elkouny. New DES based on elliptic curves. Journal of Discrete Mathematical Sciences and Cryptography, 14(4):305–315, August 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=04.

Amutha:2011:PAA

[454] R. Amutha. Performance analysis of 64-ary triangular quadrature amplitude modulation. Journal of Discrete Mathematical Sciences and Cryptography, 14(4):317–332, August 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=04.

Stoddart:2011:GHH

[455] D. R. Stoddart. The groups H3 and H4 of a class of certain cyclically presented groups are 2-generated. Journal of Discrete Mathematical Sciences and Cryptography, 14(4):333–340, August 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=04.

Tang:2011:PES

[456] H. C. Tang, K. H. Hsieh, and C. J. Wang. A partial exhaustive search for good two-term third-order multiple recursive random number generators. Journal of Discrete Mathematical Sciences and Cryptography, 14(4):341–348, August 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=04.

Karthikeyan:2011:MCT

[457] T. Karthikeyan, S. J. Peter, and S. Chidambaranathan. Meta clusters through minimum spanning tree based clustering for performance analysis of students. Journal of Discrete Mathematical Sciences and Cryptography, 14(4):349–367, August 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=04.

Das:2011:UCA

[458] S. Das and D. Roy Chowdhury. On usage of cellular automata in strengthening stream ciphers. Journal of Discrete Mathematical Sciences and Cryptography, 14(4):369–390, August 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=04.

Vijayarangan:2011:CPB

[459] N. Vijayarangan. Communication protocols based on CAAA attributes. Journal of Discrete Mathematical Sciences and Cryptography, 14(4): 391–404, August 2011. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=14&issue_id=04.

Anonymous:2011:CVAb

[460] Anonymous. Contents — Vol. 14 # 04 August 2011. Journal of Discrete Mathematical Sciences and Cryptography, 14(4):A, August 2011. CO-

DEN ???? ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=14&issue_id=04.

Peter:2011:NDM

[461] S. J. Peter. A novel dynamic minimum spanning tree based clustering method for image mining. Journal of Discrete Mathematical Sciences and Cryptography, 14(5):405–419, October 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=05.

Arumugam:2011:DDS

[462] S. Arumugam and L. Martin. Degrees and degree sequence of kedge d-critical graphs. Journal of Discrete Mathematical Sciences and Cryptography, 14(5):421–429, October 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=05.

Dayantis:2011:GAN

[463] J. Dayantis. Generalised Arnold numbers and series. Journal of Discrete Mathematical Sciences and Cryptography, 14(5):431–444, October 2011.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=14&issue_id=05.

Eteme:2011:MVA

[464] F. A. Eteme and J. A. Tsimi. A modal [. .]-valent approach of the notion of code. Journal of Discrete Mathematical Sciences and Cryptography, 14(5):445–473, October 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=05.

Peter:2011:DOH

[465] S. J. Peter. Detection of outliers and hubs using minimum spanning tree based on analytical perspective of degree numbers. Journal of Discrete Mathematical Sciences and Cryptography, 14(5):475–488, October 2011.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=14&issue_id=05.

Peter:2011:NSA

[466] S. J. Peter. Nonparametric search algorithm for outliers and hubs using minimum spanning tree. Journal of Discrete Mathematical Sciences and Cryptography, 14(5):489–502, October 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=05.

Anonymous:2011:CVO

[467] Anonymous. Contents — Vol. 14 # 05 October 2011. Journal of Discrete Mathematical Sciences and Cryptography, 14(5):A, October 2011.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=14&issue_id=05.

Sharma:2011:GKP

[468] B. D. Sharma and N. Sookoo. Generalized Krawtchouk polynomials and the complete weight enumerator of the dual code. Journal of Discrete Mathematical Sciences and Cryptography, 14(6):503–514, December 2011.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=14&issue_id=06.

Antonopoulou:2011:IDS

[469] H. Antonopoulou, N. Glinos, and Y. C. Stamatiou. An identity derived from the solution of a class of differential equations for the evolution of a key agreement protocol. Journal of Discrete Mathematical Sciences and Cryptography, 14(6):515–520, December 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=06.

Ferrero:2011:PDH

[470] D. Ferrero, S. Varghese, and A. Vijayakumar. Power domination in honeycomb networks. Journal of Discrete Mathematical Sciences and Cryptography, 14(6):521–529, December 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=06.

Li:2011:NAD

[471] R. Li. A note on the average distance of a graph. Journal of Discrete Mathematical Sciences and Cryptography, 14(6):531–536, December 2011.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=14&issue_id=06.

Chen:2011:RBR

[472] C. K. Chen. Relation between the releasing strategy and the domestic rentals on the U.S. Major Film. Journal of Discrete Mathematical Sciences and Cryptography, 14(6):537–558, December 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=06.

Eteme:2011:MAA

[473] Fidéle Ayissi Etémé and Jean Arnaud Tsimi. A mΘ approach of the algebraic theory of linear codes. Journal of Discrete Mathematical Sciences and Cryptography, 14(6):559–581, December 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=06.

Peter:2011:EDA

[474] S. J. Peter and S. Chidambaranathan. An efficient divisive-agglomerative hierarchical clustering algorithm using minimum spanning tree. Journal of Discrete Mathematical Sciences and Cryptography, 14(6):583– 595, December 2011. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-003072&
volume=14&issue_id=06.

Zhao:2011:NSC

[475] K. Zhao, Z. Li, and D. Chen. New sufficient condition and Hamiltonian and traceable. Journal of Discrete Mathematical Sciences and Cryptography, 14(6):597–602, December 2011. CODEN ???? ISSN

0972-0529. URL http://www.connectjournals.com/achivestoc.php?
bookmark=CJ-003072&volume=14&issue_id=06.

Anonymous:2011:IV

[476] Anonymous. Index of Volume 14 (2011). Journal of Discrete Mathematical Sciences and Cryptography, 14(6):603–606, December 2011. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=14&issue_id=06.

Anonymous:2011:CVD

[477] Anonymous. Contents — Vol. 14 # 06 December 2011. Journal of Discrete Mathematical Sciences and Cryptography, 14(6):A, December 2011.

CODEN ???? ISSN 0972-0529. URL http://www.connectjournals.
com/achivestoc.php?bookmark=CJ-003072&volume=14&issue_id=06.

Dass:2000:DMA

[478] B. K. Dass, editor. Discrete mathematics and allied topics. Academic Forum, Delhi, India, 2000. ISSN 0972-0529. i–iv and 1–262 pp. J. Discrete Math. Sci. Cryptography 3 (2000), no. 1–3.

Corsini:2003:AHA

[479] P. Corsini, B. K. Dass, M. Gionfriddo, and R. Migliorato, editors. Algebraic hyperstructures and applications. Taru Publications, New Delhi, India, 2003. ISSN 0972-0529. i–vi and 121–256 pp. J. Discrete Math. Sci. Cryptogr. 6 (2003), no. 2-3.